### v2

# The Impact of Surface Co-location and Eye-tracking on Mixed Reality Typing

Abstract.
Accuracy and speed are pivotal when typing. We hypothesized that the lack of tactile feedback on midair mixed reality keyboards may adversely impact performance. Our first experiment investigated providing tactile feedback to users typing in mixed reality by co-locating the virtual keyboard on a table or a wall. The keyboard was deterministic (without auto-correct), allowed mixed case typing with symbols, and relied only on the hand-tracking provided by a commodity headset’s egocentric cameras. Users preferred and had the highest entry rate of 12 words-per-minute using a midair keyboard. Error rates were similar in all conditions. Based on user feedback, our second experiment explored ten-finger typing. We used a novel eye-tracking technique to mitigate accidental key presses. This technique was preferred for ten-finger typing and halved the number of times backspace was pressed. However, participants were faster using only their index fingers without eye-tracking at 11 words-per-minute.
1. Introduction
Mixed reality (MR) offers the ability to project virtual objects onto the real world and interact with them. As MR technology develops, it may begin to displace other portable computing devices such as cell phones and laptops. Users might wish to utilize MR to perform everyday tasks. Typing is critical to efficiently performing common tasks such as writing emails, browsing the web, and creating documents. Unlike a traditional keyboard where you can feel the locations of the keys and when they are actuated, MR systems often employ midair keyboards which lack tactile feedback. The loss of tactile feedback can impair users’ accuracy when typing (Rabin and Gordon, 2004), potentially reducing the utility of MR compared to conventional devices for supporting common, everyday computing activities.
Speech input offers an alternative to typing but may not always be socially acceptable (Baier and Burmester, 2019), and recognition of difficult words can be a problem (Adhikary and Vertanen, 2021). Auxiliary devices such as physical keyboards (McGill et al., 2015; Walker et al., 2017; Knierim et al., 2018; Grubert et al., 2018), controllers (Speicher et al., 2018a; Xu et al., 2019a), gloves (Markussen et al., 2014; McCaul and Sutherland, 2004; He et al., 2022), or watches (Ahn et al., 2017), could help but require carrying additional devices. Other solutions re-imagine the text input process (e.g. word-gesture keyboards (Kristensson and Zhai, 2004; Zhai and Kristensson, 2012; Ahn et al., 2017), dwell-free gaze typing (Kristensson and Vertanen, 2012; Hu et al., 2024), or finger gestures (McCaul and Sutherland, 2004; He et al., 2022; Feit et al., 2015)). However, such methods introduce a learning curve that may be a barrier to MR adoption. Compared to past work, we focus on enabling mixed reality typing without: 1) requiring users to learn a new interaction technique, 2) requiring devices or sensors aside from the headset, and 3) using a language model to compensate for noisy input.
We conjectured providing tactile feedback would improve typing in MR. In Experiment 1, we aligned a virtual keyboard with a wall or table. This adds tactile feedback to keypresses and could be used whenever a flat surface is present. Although surface co-location may not be feasible in all situations, people are frequently in human-made environments where flat surfaces are prevalent. We compared the surface aligned keyboards with one floating in midair. Participants typed using only their index fingers. We found participants preferred the midair keyboard compared to keyboards near a table or wall. They typed faster in midair at 12.2 words-per-minute (WPM) compared to on a table (9.1 WPM) or a wall (8.8 WPM). Experiment 2 focused on ten-finger typing, a feature requested by many participants in our first study. Pilot testing revealed ten-finger typing in midair led to numerous accidental keystrokes. Therefore, in addition to comparing ten- and index-finger typing, we also added a novel eye-tracking feature designed to prevent accidental keystrokes. Our feature filtered out key presses not located near a user’s eye gaze location. We theorized since users were likely visually guiding their typing due to the lack of tactile feedback, this feature would not be an undue burden. We found users preferred index- to ten-finger typing. We found eye-tracking reduced accidental key presses, improved accuracy, and slightly increased entry rates for ten-finger typing (7.4 WPM versus 7.1 WPM).
Our contributions are as follows:
In two studies with 42 total participants, we found users could enter rich text (i.e. mixed case with punctuation and numbers) accurately and at a reasonable speed of 11–12 words-per-minute on a deterministic midair keyboard.
User feedback was in general positive for the tactile feedback and ergonomics offered by virtual keyboards located near a surface. However near surfaces, accurate hand-tracking was a challenge for a current commercial headset.
Our novel eye-tracking based filter feature reduced unintentional key presses in midair ten-finger typing.
We release detailed logs from our studies and the source code of our HoloLens keyboard experimental applications.
2. Related Work
If mixed reality is going to be used to perform everyday computing tasks, text input needs to be efficient. Many solutions have been proposed to address this need. A midair virtual keyboard that allows users to type directly on it with their hands is both familiar and easy-to-learn. Early systems such as ARKB (Lee and Woo, 2003) relied on external cameras to track a user’s hands over a virtual keyboard. Markussen et al. (Markussen et al., 2013) used an array of 24 external cameras to track a user’s hands. Users directly selected keys on a projected QWERTY keyboard, or used ambiguous gestures to specify keys. Direct typing was the fastest at 13 WPM after six sessions. Dudley et al. (Dudley et al., 2018) investigated midair virtual keyboard typing on a HoloLens 1. Similar to our work, they relied only on the built-in headset sensors. Despite the HoloLens 1 only tracking a user’s wrist position, users typed at 18 WPM on an auto-correcting keyboard with word predictions. In comparison, we used a HoloLens 2 with full hand-tracking and focused on input of more difficult mixed case text with symbols and numbers.
Cheng et al. (Cheng et al., 2022) explored if sitting at a table improved users’ ability to select and drag targets in virtual reality (VR). Users’ arms and finger were tracked via external cameras. They found the presence of a table improved user accuracy and reduced fatigue of horizontal and vertical midair interfaces. In this work, we also chose to use a table as it is a common everyday object. However, we focus on MR and the more accuracy demanding task of keyboard typing.
While keyboard typing is a familiar interaction, the lack of tactile feedback on a midair keyboard can be a challenge. Tactile feedback significantly aids users even when typing on a physical keyboard. Typists with anesthetized fingertips were found to make seven times as many errors (Rabin and Gordon, 2004). Tactile feedback can be added to these virtual keyboards via a handheld controller (Boletsis and Kongsvik, 2019; Speicher et al., 2018b), a physical keyboard (McGill et al., 2015; Walker et al., 2017; Knierim et al., 2018; Grubert et al., 2018), or devices placed on the fingers, hand, or wrist (Gupta et al., 2020; Pezent et al., 2019; Sakashita et al., 2019; Gu et al., 2016). But this requires users carry these devices in addition to the headset. Utilizing existing surfaces to provide tactile feedback avoids the needs for additional devices and has shown promising results. Users can type well on flat surfaces (Findlater et al., 2011) such as a tabletop display and on virtual keyboards projected onto a flat surface (Roeber et al., 2003; Dudley et al., 2019).
Dudley et al. (Dudley et al., 2019) had users type with two or ten fingers on virtual keyboards located near a surface or in midair. Users’ hands were tracked via an array of external cameras and markers on the hands. The system used an auto-correct algorithm based on the system knowing the target text. Users typed significantly faster on the surface keyboards with either two fingers (56 WPM versus 42 WPM) or all ten fingers (52 WPM versus 35 WPM). Richardson et al. (Richardson et al., 2020) also investigated surface typing by tracking users’ hands via external cameras. Users typed phrases quickly at 73 WPM. Using a neural network hand motion model and a language model, they found recorded typing sequences could be decoded at a low error rate of 2%. In recent followup work, Richardson et al. (Richardson et al., 2024) used only the egocentric cameras of a VR headset to infer keypresses on a virtual keyboard situated on a desk. Users typed at 44 WPM with an uncorrected error rate of 7%. While the system had no explicit language model, the motion model may have learned one as the majority of its training data was English text. Indeed, on non-language text, offline experiments showed a much high error rate of 16% compared to 5% on English. Similar to (Richardson et al., 2024), we utilize the hand-tracking available from the egocentric cameras of a commodity headset. However, we rely solely on a human-in-the-loop process where the user observes their inferred finger locations and uses this feedback to supervise their typing. A system free from language-based prediction bias may be needed in some scenarios (e.g. typing a password).
Moving somewhat away from the metaphor of tapping on a keyboard, a word-gesture keyboard (Kristensson and Zhai, 2004; Zhai and Kristensson, 2012) allows writing by swiping through a word’s letters. The Vulture keyboard (Markussen et al., 2014) allowed writing on a keyboard using a glove and an external camera array. After ten sessions, users wrote at 21 WPM. Moving further away from conventional typing, an algorithm can infer a user’s intended text from ambiguous gestures. Early work allowed users wearing a glove to write words in VR by selecting different sets of letters by flexing their finger (McCaul and Sutherland, 2004). A similar recent system is TapGazer (He et al., 2022) which uses a glove to detect which finger a user is tapping with disambiguation via eye gaze or additional taps.
Other approaches for text input in VR and AR have explored input methods other than hand gestures. The direction of a user’s head can be used to select a letter from a circular menu (Xu et al., 2019b) or a QWERTY keyboard (Yu et al., 2017; Speicher et al., 2018a; Gizatdinova et al., 2018; Xu et al., 2019). While Gizatdinova et al. (Gizatdinova et al., 2018) found head-based typing was effective on small keyboards, users thought it felt unnatural and they preferred gaze-based typing. Gaze-based typing in VR has been investigated both as the sole input method (Lu et al., 2021) and in combination with head gestures (Feng et al., 2021). However, the use of dwell time, the minimum time a user must look at a key in order to trigger it, can lead to eyestrain (Lystbæk et al., 2022; Pfeuffer et al., 2020). An alternate approach, dwell-free eye typing (Kristensson and Vertanen, 2012), allows users to gaze continually through letters. Using this approach, Hu et al. (Hu et al., 2024) found users wrote at 12 WPM using a HoloLens 2 headset. While potentially fast, dwell-free typing requires user learning and may not work for hard-to-predict text.
Using speech input can help reduce MR challenges such as robustly sensing users hands and accurately displaying visual feedback. In early work, Bowman et al. (Bowman et al., 2002) explored VR text input using Wizard of Oz speech recognition and input via a glove, keyboard, or tablet. Speech recognition was fastest at around 13 WPM. In SWIFTER (Pick et al., 2016), users wrote in a virtual environment at 24 WPM using speech with correction via a hand controller. Similar to our system, Adhikary and Vertanen (Adhikary and Vertanen, 2021) eschewed the use of auxiliary devices and relied solely on speech and hand-tracking for VR text input. Speaking and then correcting via hand gestures was the fastest at 28 WPM. Notably, without speech, users wrote at 11 WPM on a midair auto-correcting keyboard with word predictions. This entry rate was similar to our users’ midair deterministic keyboard performance. While speech input can be a good option, it may not always be socially acceptable (Baier and Burmester, 2019), and accuracy can be poor in noisy environments or for challenging text (e.g. usernames and passwords).
Most existing MR typing solutions, while often allowing typing with both hands, typically only support typing with the index fingers (Lystbæk et al., 2022; Dudley et al., 2018; Adhikary and Vertanen, 2021). Systems that do allow use of all fingers generally require an auxiliary input device (He et al., 2022), external sensing device (Yi et al., 2015), or physical keyboard (McGill et al., 2015; Walker et al., 2017; Knierim et al., 2018; Grubert et al., 2018). Ten-finger typing is challenging as the dexterity limitations of the human hand can make some multi-finger gestures difficult to perform without the tactile assistance of a physical keyboard (Sridhar et al., 2015). Additionally, ten-finger typing on virtual keyboards can lead to accidental inputs and key coactivations (Foy et al., 2021). While multi-finger gestures have been used for navigating MR interfaces, the small movements required for typing may be difficult for vision-based hand-tracking systems to recognize (Malik et al., 2005). To try and improve ten-finger typing, in Experiment 2, we use eye-tracking to filter out key presses not located near a user’s gaze. Past work shows that users will naturally tend to look at the keys they are pressing (Lystbæk et al., 2022).
3. Experiment 1
In our first experiment, participants wore a HoloLens 2 headset and typed sentences on a virtual keyboard using their index fingers. We measured user performance and preference for typing on a midair keyboard and keyboards co-located with a table and with a wall. Our primary goal was to assess the impact of the tactile feedback offered by the surface co-location. We opted not to include a physical keyboard condition since typing on physical keyboards has been well-studied and our focus is on on-the-go applications where all sensing is performed by the headset. Since midair typing is currently a common choice for MR typing applications, we used a midair keyboard as a baseline.
3.1. Design
The experiment consisted of three conditions (depicted in Figure 1):
Table — The keyboard was placed just above a table, using a QR code on the table to ensure consistent spatial placement. The keyboard was placed horizontally, such that the back of the keyboard was parallel with the table.
Wall — The keyboard was placed slightly in front of a wall, using the same QR code as a spatial anchor. The keyboard was placed vertically, such that the back of the keyboard was parallel with the wall.
We hypothesized the tactile feedback conditions (i.e. Wall and Table) would perform better than the non-tactile feedback condition (i.e. Midair). We also hypothesized Table would be preferred due to its similarity to a standard physical keyboard. We investigated both the Wall and Table locations to assess typing on common real-world flat surfaces. In addition, the keyboard position and orientation relative to the headset is different between the Wall and Table conditions. This allowed us to investigate whether this impacted hand-tracking accuracy.
In all conditions, participants used their index fingers on one or both hands to enter text. This choice prioritizes ecological validity by allowing participants to type as they would in the real world. Participants also could choose to rest their arms on a table as needed to potentially reduce fatigue. Our keyboard was deterministic and did not auto-correct or predict words. We did this to understand the raw performance of typing as might be required for text not easily predicted by a language model.
All conditions had audio and visual feedback for key presses. A key was triggered when either index finger entered the key’s bounding area, pushed the key’s face a certain depth, and exited the same key’s bounding area. On a key press, we played a button sound effect and displayed a white border around the key. We displayed participants’ hands using the Unity MR toolkit’s default hand skin, an overlay of white and grey triangles. We displayed blue spheres on fingertips and grey cubes on hand joints (Figure 1 right). The keyboard was 44 cm 16 cm and rendered slightly transparent blue with white key labels. We used a full QWERTY keyboard including numbers, symbols, shift keys, caps lock, and backspace. Users could see their current text above the keyboard and could correct errors via a backspace key.
A small table with an office chair, a keyboard floating midair in front of the user, a keyboard placed on the horizontal table surface, and a keyboard placed on the wall behind the table.
3.2. Procedure
Participants used a HoloLens 2 headset while seated at a small table (Figure 1 left). We told participants they could type using the index fingers on one or both hands. During the calibration phase, participants used only one hand to avoid potential depth measurement issues caused by sensing the non-typing hand. During the practice and evaluation sentences, we reminded users they could type with both hands. We allowed participants the freedom to type with one or both hands as it better simulates real-world keyboard typing where users have the same freedom.
The study lasted approximately one hour. Participants completed an initial questionnaire before starting the conditions. Between conditions, participants took a two-minute break. During the break, we instructed participants to lift the HoloLens visor to reduce eyestrain. They then completed a questionnaire about the previous condition. After all conditions, participants completed a final questionnaire. On the questionnaires, participants ranked various statements on a 7-point Likert scale (1=strongly disagree, 7=strongly agree). We also asked for open responses to various questions. For our exact questionnaires, see our supplementary materials.111https://osf.io/q4n72 Participants received a $15 USD gift card.
The order of conditions was completely counterbalanced between participants. Prior to each condition, participants typed a calibration sentence to ensure the keyboard’s consistent placement relative to the surface. If the keyboard was too far in front of the surface, the finger could actuate a key without tactile feedback. If it was too close to the surface, the tracked finger would hit the real surface before it actuated a virtual key—making it impossible to type. Calibration computed the average depth of nine key presses when the participants typed “calibrate”. The participants were instructed to keep pushing the key down until they felt the surface. While the calibration sentence only affected the Wall and Table conditions, we retained it in the Midair condition for consistency.
Following calibration, participants typed two practice sentences and twelve evaluation sentences. We only analyzed the evaluation sentences. Sentences were pulled randomly without replacement from the “mem1-5” set of the mobile Enron dataset (Vertanen and Kristensson, 2011). These sentences were in mixed case and included punctuation and sometimes numbers.
Eighteen participants between the ages of 18 and 26 took part in Experiment 1. 15 participants identified as male, 2 as female, and 1 as non-binary. 16 participants strongly agreed with the statement “I consider myself a fluent speaker of English”, one participant agreed, and one participant strongly disagreed. We believe the participant who strongly disagreed had made an error when marking the scale. Eight participants reported augmented reality (AR) experience, including one participant who reported owning a HoloLens. The remaining seven participants reported limited AR experience, primarily with phone apps such as Pokémon Go.
Throughout this paper, we test for statistical differences using repeated measures ANOVA. In cases were Mauchly’s test indicated a violation of sphericity, we report Greenhouse-Geisser corrected tests. For significant omnibus tests, we test for differences between conditions via Bonferroni corrected post-hoc tests. For Likert ratings, we used Friedman’s ANOVA. For significant omnibus tests, we test for differences between conditions via a Bonferroni corrected Wilcoxon signed-rank test.
3.3. Results
Entry rate was measured in words-per-minute (WPM). We considered a word to be five characters (including space). We subtracted one from the number of characters to compensate for the unmeasured time finding and typing the first key (Wobbrock, 2007). We started timing from when the first key was pressed. We ended timing when the next button was pressed since this includes the perceptual, cognitive, and motor overheads associated with confirming the last character typed. Entry rates were 12.2, 9.1, and 7.8 WPM in the Midair, Table, and Wall conditions respectively (Figure 2 left). This difference was significant (Table 1 left). Post-hoc tests found Midair was significantly faster than both Table and Wall.
Entry rate (WPM) | Error rate (CER %) | Backspaces per final character | |
Midair | 12.2 2.1 [8.8, 17.6] | 0.95 1.9 [0.0, 7.8] | 0.119 0.092 [0.017, 0.330] |
Table | 9.1 2.2 [5.8, 13.1] | 0.98 2.1 [0.0, 9.2] | 0.160 0.151 [0.015, 0.612] |
Wall | 7.8 1.3 [5.9, 11.2] | 0.65 1.2 [0.0, 4.4] | 0.132 0.091 [0.040, 0.349] |
Omnibus test | = , , | , , | , , |
Pairwise | Midair Table ( ¡ ) | n/a | n/a |
post-hoc | Midair Wall ( ¡ ) | ||
Table Wall ( = ) |
We measured typing errors using character error rate (CER). CER refers to the number of character insertions, substitutions, and deletions required to change a participant’s final entered text to the reference text, relative to the characters in the reference and multiplied by 100. Error rates were low in all conditions with a CER of 0.95%, 0.98%, and 0.65% in Midair, Table, and Wall respectively (Figure 2 center). This difference was not significant (Table 1 center).
As a measure of how much effort participants were expending correcting errors, we calculated the number of backspaces per final output character for each phrase. While the final character error rates in the conditions were relatively low, we observed frequent use of backspace. The number of backspaces pressed per output character was 0.12 for Midair, 0.16 for Table, and 0.13 for Wall. While there did seem to be more backspacing in the two surface co-located keyboards (in particular for Table), this difference was not significant (Table 1 right).
Three box plots showing the entry rate, character error rate, and backspaces per character in Experiment 1. Each plot shows the results for each of the three conditions.
Scatter plot showing how each participant did in each condition of Experiment 1 with respect to the character error rate (x-axis) and entry rate (y-axis).
We were curious how performance differed among participants. Figure 3 shows each participant’s performance in each condition. In Midair, except for two participants, everyone wrote at 10 WPM or faster with a CER of less than 3%. In all conditions, with the exception of two participants, everyone obtained a final error rate of less than 5%. In all conditions, many participants corrected all errors to achieve exactly the requested text (including capitalization and punctuation). This demonstrates typing on a virtual keyboard is a viable, albeit somewhat slow, mixed reality text entry method. This includes for text that cannot be auto-corrected (e.g. passwords, account numbers, and URLs).
We recorded the position of participants’ left and right index fingers 30 times per second. We also noted whenever tracking was lost for either finger. During the time periods participants were typing the evaluation phrases, we computed the percentage of a participant’s samples in which at least one finger was tracked. We found tracking success was 99.9%, 99.1%, and 98.6% in Midair, Table, and Wall respectively. While this difference was significant ( = , , ), post-hoc pairwise comparisons were not significant: Midair Table ( = ), Midair Wall ( = ), and Table Wall ( = ). We note that just because we had a tracked location of the index fingers, this does not mean that location was necessarily accurate. Subjective feedback from many participants indicated hand-tracking was less accurate in Midair and Table. We conjecture this could have resulted from a number of factors such as the proximity of the hands to the surface, occlusion of the fingers or hands between themselves, or changes due to the relative location and angle between the hands and the headset.
Three figures of a QWERTY keyboard. Each letter is labeled with the percentage of attempts to type that letter that were successful.
To investigate if participants’ accuracy was impacted by the spatial location of their desired key, we calculated the percentage of time participants successfully obtained the next letter in the target phrase. We calculated this only on key presses in which, at the time of the key press, a participant’s current typing result was a strict prefix of the target. For example, if the target phrase was “Are you there?” and the participant had typed “Are you t” we would score the next key press based on a target of “h”. However if they had typed “Are you h” we would drop the next key press as we are unsure if they accidentally missed the “t” key or if they are writing an incorrect word like “here”. Filtering in this way allows us to be reasonably confident about a participant’s next intended key. As shown in Figure 4, there did not seem to be a strong difference in accuracy based on spatial location in Midair or Table. However in the Wall condition, the bottom keyboard row was noticeably less accurate at 85% versus 91% for the middle row and 92% for the top row.
3.4. Subjective feedback
Image showing the Likert ratings for three statements in Experiment 1.
After each condition, participants rated three statements on a 7-point Likert scale (Figure 5). Participants’ average ratings for “I entered text quickly” were: Table 4.0, Wall 2.9, and Midair 5.1. The difference between ratings was significant (). The only significant post-hoc test was that Midair was quicker than Wall (). Participants’ average ratings for “I entered text accurately” were: Table 4.5, Wall 3.8, and Midair 5.4. The difference between ratings was significant (). The only significant post-hoc test was that Midair was more accurate than Wall (). Participants’ average ratings for “Entering text was physically easy and not tiring” were: Table 4.3 for, Wall 2.9, and Midair 4.4. The difference between ratings was significant (). The only significant post-hoc test was that Midair was physically easier than Wall ().
After each condition, we asked participants what they most liked and most disliked about the condition. Table 2 shows selected comments about each condition. Overall, a common like in Table and Wall was the tactile feedback afforded by the physical surface. However, in both conditions, a common dislike was problems related to the tracking of hands and failure to register key presses. Participants complained of physical strain in all conditions, though Table received the most positive feedback regarding ergonomics. For all participant comments, see our supplementary materials.222https://osf.io/q4n72
Condition | Comment | |
Midair | + | “Easier to lean arms against table.” |
+ | “Not needing to touch the surface behind me helped me to lift my fingers up farther, reducing misinput.” | |
+ | “My hands wanted to be there naturally and didn’t require as much movement between letters.” | |
- | “Because my hands were in the air, it made physically more exhausting for longer period of time.” | |
- | “There’s no tactile feedback.” | |
- | “Without the wall or table it was difficult to estimate the depth of the keyboard.” | |
Table | + | “I got to keep my arms on the table.” |
+ | “It was a more natural typing posture.” | |
+ | “Having the table under the keyboard allowed for a form of tactile response.” | |
- | “It seemed I lost tracking a lot in this part; the 3D outlines of the keys felt disorienting.” | |
- | “Accurate input required me to fully focus on my hands, causing back strain.” | |
- | “I had to look further down than normal.” | |
Wall | + | “I liked that the keyboard would type pretty much right as I hit the wall.” |
+ | “Using this interface, I tended to forget that the keyboard wasn’t physically in front of me.” | |
+ | “I didn’t have to pull away so much between letters. I didn’t really touch the table but touched the wall more.” | |
- | “Awkward position, tiring.” | |
- | “Hard to visualize where I needed to put my hands to type best. Typing felt very finicky.” | |
- | “I felt like the keyboard strokes were more inaccurate than the other two tests.” |
In a final questionnaire, participants were asked to rank the conditions. Of the 18 participants, 13 preferred Midair, 5 preferred Table, and none preferred Wall. We asked participants to explain their ranking of the conditions. Of those who ranked Midair the best, some representative comments were: “midair was the fastest and tracked most accurately”, and “hovering was most comfortable and easy to use”. Of those ranking Table as the best, participants commented: “horizontal was natural feeling and responsive”, and “horizontal provided the most comfort and accuracy for me”. Most participants ranked Wall as the worst and participants often mentioned physical strain: “vertical was the slowest and most physically taxing”, and “wall was neither comfortable nor accurate”.
Finally, we asked participants “what needs to be improved or changed about the AR headset’s hardware or software before you would be willing to use it in place of your mobile phone?” Requested changes were:
Improved hand-tracking accuracy and responsiveness (10 participants).
Interaction with multiple fingers (10 participants).
Improvements related to the visual capabilities of the headset (6 participants).
Auto-correct or word predictions (3 participants).
Ability to locate the keyboard wherever you wanted it (1 participant).
Swipe (word-gesture keyboard) typing (1 participant).
Smaller form factor headset (1 participant).
3.5. Discussion
Our goal in Experiment 1 was to assess whether typing with tactile feedback outperformed or was preferred compared to typing in midair. We conjectured that the tactile feedback conditions, especially the Table condition which is similar to a standard physical keyboard, would perform better and be more preferred. However, our data showed the Midair condition had a significantly higher entry rate than both Table and Wall. All conditions had low final error rates, though moderate use of backspace was observed. Furthermore, in the final questionnaire, thirteen of the eighteen participants preferred the Midair condition, with the second most preferred condition being the Table condition by the remaining five people. We observed throughout the study that the Table and Wall conditions reduced the reliability of key actuation and this was corroborated by comments made by participants. This may explain the increased performance and preference for the Midair condition.
We believe the poor performance in the Table condition may be due to our calibration process or the hand-tracking. During calibration, the participant had to type the word “calibrate”. We instructed participants to continue pushing the key until their finger touched the surface. This form of calibration was dependent on how precise the participant was. We found that with quick hand movements, the HoloLens’ hand-tracking lagged and when the hand abruptly stopped, the resulting depth measurement could be deeper than the participant’s finger had actually gone. The latency of the hand-tracking may also have caused inaccuracies, Kim et al. (Kim et al., 2023) report a latency of 90 ms and tracking jitter of 1 mm using a HoloLens 2 headset. It might be possible to improve the keyboard’s placement above the surface with a modified or extended calibration procedure. It might also be advisable to allow users to interactively adjust the keyboard height during typing. A different tracking system may have produced different results. However, this work provides valuable information about how well surface co-location works given the tracking of a commonly available MR headset.
We also speculate the unreliable key triggering is related to limitations of the HoloLens’ egocentric cameras or its hand-tracking models. During the experiment, we saw participants try to compensate for problems with hand-tracking in a number of ways, such as moving closer to the keyboard or adjusting the viewing angle between the headset and the keyboard. However, in some cases, such adjustments only made tracking less reliable. Ideally, the HoloLens interface would provide some sort of informative feedback that could guide users to adjust their position or environment to improve tracking. Presently, the only feedback is the unstable rendering of a user’s hand skeleton.
Midair interaction can be tiring since users have to hold their hands up with little support (Boring et al., 2009; Hincapié-Ramos et al., 2014). Given this, we were surprised to find participants rated Midair the least tiring (Figure 5 bottom). As we expected, the Table was also comfortable, likely because people could rest their arms on the table similar to traditional keyboard typing. The Wall condition was reported to be the most physically demanding. Although participants could rest their elbows on the table in this condition, it might have worked better for some participants than others. This condition is also the most different from traditional typing. For consistency with the other conditions, participants stayed seated at the table while typing on the wall. It could be that wall typing would be more ergonomic while standing.
4. Experiment 2
Experiment 1 showed a significantly higher entry rate for the midair keyboard and users strongly preferred that keyboard. The majority of participants also requested the ability to type with multiple fingers. Based on this, we decided to focus exclusively on a midair keyboard in Experiment 2. Our goals were to 1) explore if ten-finger typing is feasible on a midair keyboard, and 2) determine if eye-tracking could aid in eliminating accidental key presses.
4.1. Design
In pilot testing of ten-finger typing, we found users had difficulty maintaining a consistent hand position in midair. Monitoring the positions of all ten fingers was difficult and this resulted in a high number of accidental key presses. The lack of tactile feedback also meant users often did not realize they had unintentionally pressed a key. Because of this, we worried that ten-finger typing, on its own, would not be competitive with index finger typing.
Increasingly, VR and AR headsets (including the HoloLens 2) have built-in eye-tracking. We hypothesized users would likely be looking at or near the key that they are trying to press. By tracking users’ eye gaze locations, we could filter out key presses that were sufficiently far from where they are looking. We hoped that this would improve efficiency by reducing the time spent carefully monitoring hand positions and correcting accidental key presses.
Like Experiment 1, the keyboard was deterministic without auto-correct or word predictions. In the conditions using the eye-tracking filtering, we enabled all keys within a 2.6 cm radius of the user’s gaze location (roughly a one key radius). Keys outside this radius were disabled. Disabled keys were gray, while enabled keys were transparent (Figure 7). When pressed, a key flashed a slightly transparent blue and a sound was played. All keyboard text was white.
Experiment 2 consisted of four conditions:
Index — We asked participants to type only with their index fingers. Key presses from other fingers were ignored.
IndexEye — We asked participants to type only with their index fingers and to look at the keys as they typed. Key presses from other fingers were ignored as were keys outside of the radius around their gaze location.
Ten — We asked participants to type with all ten fingers. Key presses were triggered by all fingers.
TenEye — We asked participants to type with all ten fingers and to look at the keys as they typed. Key presses were registered for all fingers. Keys outside of the radius around their gaze location were ignored.
The first image shows the interface used in Experiment 2. It consists of a keyboard, a middle area where the user’s typed text appears, and an upper area that displays the phrase the user is copying. The second image shows a user typing with the index fingers of both hands. Each index fingertip has a red sphere on it. The third image shows a user typing with all ten fingers. Each fingertip has a red sphere on it.
Image of a full QWERTY keyboard. All the keys have a grey background except for the H key and the keys adjacent to it which all have a transparent background.
A key was triggered when a participant’s finger pushed a key downward past a threshold of approximately one-third of a centimeter. In the event that multiple keys were pressed at the same time, only the furthest pressed key was registered as pressed. Following a key press, there would be a half-second cool down where the keyboard would not register any key presses to help prevent multiple keys from being triggered from a single tap gesture.
We used a QWERTY keyboard including numbers, symbols, shift keys, caps lock, and backspace. The keyboard was 35 cm 12 cm, with each letter key being 2.2 cm 2.2 cm. Red spherical colliders were attached to the user’s index fingers in Index and IndexEye, and all ten fingers in Ten and TenEye.
We opted for a visual design mimicking a laptop (Figure 6) since there was no need for the interface to appear in a single plane (as was the case to support the Wall condition in Experiment 1). The top virtual screen shows the phrase the participant is being asked to type. The center text area shows the current text typed thus far. Typing occurs on the lower keyboard area, with the end of a sentence signaled via the large enter key on the right. A backspace key allows deletion of previously entered characters. As in Experiment 1, the keyboard was tilted 15 degrees above horizontal.
4.2. Procedure
Participants used a HoloLens 2 headset while seated. The keyboard was placed midair in front of the participant. Participants were instructed to adjust the height of the chair such that they could comfortably reach the keys. Since the keyboard was intended to be used in midair only, no table was included in Experiment 2.
The study took approximately one hour. Participants first completed an initial questionnaire and then proceeded to their first condition. Between conditions, participants took a two-minute break and completed a questionnaire about the condition just finished. During this time, we instructed them to remove the headset and walk around the room to reduce eyestrain. After all conditions, participants completed a final questionnaire. The questionnaires were identical to Experiment 1 and are included in our supplementary materials.333https://osf.io/q4n72 Participants received a $15 USD gift card.
The order of conditions was completely counterbalanced between participants. Each condition consisted of two practice sentences, followed by eight evaluation sentences. We only analyzed the evaluation sentences. For each condition, sentences were pulled randomly without replacement from the “mem1-5” set from the mobile Enron dataset (Vertanen and Kristensson, 2011). These sentences were in mixed case and included punctuation and sometimes numbers. Participants were instructed to type using both hands.
24 participants between the ages of 18 and 22 completed Experiment 2. Of these, 17 were male, 5 were female, and 2 preferred not to say. 23 participants strongly agreed with “I consider myself a fluent speaker of English”, and the remaining participant agreed with the statement. A total of seven participants reported AR experience, with four participants having used a HoloLens before, and three participants reporting use of phone AR apps (e.g. Pokémon Go).
4.3. Results
We measured entry rate as in Experiment 1. Entry rates were 10.9 in Index, 9.4 in IndexEye, 7.1 in Ten, and 7.4 WPM in TenEye (Table 3 left and Figure 8 left). This difference was significant. All post-hoc pairwise comparisons between conditions were significant except between Ten and TenEye.
Entry rate (WPM) | Error rate (CER %) | Backspaces per final character | |
Index | 10.9 1.4 [8.5, 13.9] | 0.63 0.82 [0.00, 3.11] | 0.12 0.06 [0.01, 0.27] |
IndexEye | 9.4 1.3 [6.9, 11.5] | 0.58 0.92 [0.00, 2.89] | 0.08 0.07 [0.02, 0.31] |
Ten | 7.1 1.7 [3.1, 10.5] | 1.91 2.21 [0.00, 9.63] | 0.30 0.14 [0.07, 0.62] |
TenEye | 7.4 1.9 [4.3, 11.0] | 1.27 1.54 [0.00, 6.52] | 0.16 0.11 [0.04, 0.39] |
Omnibus test | , , | , , | , , |
Pairwise | Index IndexEye, | Index IndexEye, | Index IndexEye, |
post-hoc | Index Ten, | Index Ten, | Index Ten, |
Index TenEye, | Index TenEye, | Index TenEye, | |
IndexEye Ten, | IndexEye Ten, | IndexEye Ten, | |
IndexEye TenEye, | IndexEye TenEye, | IndexEye TenEye, | |
Ten TenEye, | Ten TenEye, | TenEye Ten, |
Three box plots showing the entry rate, character error rate, and backspaces per character in Experiment 2. Each plot shows the results for each of the four conditions.
Scatter plot showing how each participant did in each condition in Experiment 2 with respect to the character error rate (x-axis) and entry rate (y-axis).
As in Experiment 1, we measured typing errors using character error rate (CER). The CER was 0.63% for Index, 0.58% for IndexEye, 1.91% for Ten, and 1.27% for TenEye (Table 3 center and Figure 8 center). While this difference was significant, the only significant post-hoc pairwise comparison was between IndexEye and Ten.
To measure error correction activity, we calculated the number of backspaces per final output character for each phrase. The number of backspaces pressed per output character was 0.12 for Index, 0.08 for IndexEye, 0.30 for Ten, and 0.16 for TenEye (Table 3 right and Figure 8 right). This difference was significant. Post-hoc pairwise comparisons were significant with the exception of between Index and IndexEye, and between Index and TenEye. Notably, eye-tracking did significantly reduce the number of corrections required in ten-finger typing. This suggests that our filtering feature was successful at preventing some erroneous keystrokes.
As shown in Figure 9, numerous participants in Ten and TenEye had low entry rates of 4–8 WPM. Additionally, all instances of an error rate of 4% or above were participants in these two conditions. While entry rate was overall slow in the ten-finger conditions, the addition of eye-tracking did seem to shift performance towards a lower error rate.
We measured the percentage of the time users were looking at a key in the Index and Ten conditions such that the key would have been enabled in IndexEye and TenEye. This occurred for 83.8% of key presses in Index and 84.6% in Ten. Our implementation of eye-tracking eliminated 10.4% of total key presses from IndexEye and 36.7% from the TenEye. This indicates that user’s gaze and typing style was relatively consistent between the Index and IndexEye conditions, since 83.8% of key presses being valid in Index is similar to 89.6% being valid in IndexEye. There was a much larger discrepancy between the Ten and TenEye conditions, with only 63.3% of key presses being valid in TenEye as opposed to 84.6% in Ten. This may indicate that users were less careful with their hands in the TenEye condition and relied on the eye-tracking feature to filter out any extraneous key presses.
A line graph showing the average eye gaze distance of participants from a pressed key over a time offset from the key press.
Heatmap showing how often each participant used their left or right index finger.
Heatmap showing how often each participant used each of their fingers.
Figure 10 shows the average distance between the pressed key and eye gaze, before and after the key was pressed. In all conditions, the user looked most closely at the key at the instant it was pressed. Despite users tending to look more closely at the pressed key in the IndexEye and TenEye conditions, likely due to the restraints imposed by our eye-tracking feature, all four conditions followed the same pattern of eye movement. This supports our theory that users will naturally tend to look where they intend to type, and indicates potential for incorporating this implicit gaze signal to improve a keyboard’s interface or algorithms (e.g. auto-correct or word predictions).
We calculated how often participants used each of their fingers during all conditions. Figure 11 shows that participants in the Index and IndexEye conditions tended to share typing between their two hands. For example, 11 participants in the Index condition used their left index finger 40% of the time or higher. As shown in Figure 12, in the ten-finger conditions, in general participants used all their fingers. However, most participants showed a strong preference to typing with their index fingers. In the Ten condition, on average participants typed with their left and right index fingers 17% and 30% of the time respectively. Interestingly, in the TenEye condition, this tendency was less pronounced at 12% and 21% for the left and right index fingers respectively. This shows that the eye-tracking feature may have increased participants’ confidence in typing with other fingers.
4.4. Subjective feedback
After each condition, participants rated three statements on a 7-point Likert scale (Figure 13). Participants’ average ratings for “I entered text quickly” were: Index 4.4, IndexEye 3.9, Ten 3.1, and TenEye 3.3. There was a significant difference between ratings (), but none of the pairwise differences were significant. Participants’ average ratings for “I entered text accurately” were: Index 5.2, IndexEye 5.8, Ten 3.5, and TenEye 4.5. There was a significant difference between ratings (). Post-hoc tests showed Index was rated more accurate than Ten (), IndexEye was more accurate than Ten (), and IndexEye was more accurate than TenEye (). Other pairs were non-significant. Participants’ average ratings for “Entering text was physically easy and not tiring” were: Index 4.2, IndexEye 4.4, Ten 3.7, and TenEye 4.5. There was not a significant difference between ratings ().
Image showing the Likert ratings for three statements in Experiment 2.
In a final questionnaire, participants ranked the conditions and were asked to explain their choice. 16 of 24 participants ranked Index first. Of those ranking Index the best, some representative comments were: “I didn’t have to worry about other fingers getting in the way”, “index fingers are both faster and more accurate for this”, and “I prefer the comfort of being able to relax my other fingers”. The second most preferred condition, by seven participants, was IndexEye. A representative comment from a participant preferring this condition was: “index finger tracking only was the easiest, and when combined with eye-tracking, it was also the most accurate”. Only one participant preferred TenEye.
Regarding the index finger conditions, participants seemed to prefer no eye-tracking, with 17 participants ranking Index above IndexEye. When only using index fingers, most participants seemed to think that the eye-tracking in IndexEye was more limiting than helpful. In the ten-finger conditions, however, participants seemed to prefer the eye-tracking feature, with 17 participants ranking TenEye above Ten. Participant comments specific to each condition are listed in Table 4. For all user feedback, see our supplementary materials.444https://osf.io/q4n72
Finally, we asked participants “what needs to be improved or changed about the AR headset’s hardware or software before you would be willing to use it in place of your mobile phone?” Requested changes were:
Improved hand-tracking accuracy and responsiveness, particularly for the ring finger (13 participants).
Improvements related to the visual capabilities of the headset, specifically field of view (6 participants).
Reduced eye-tracking latency (5 participants).
Auto-correct or word predictions (4 participants).
Increasing the keyboard sensitivity (1 participant).
Reducing the bulk of the headset (1 participant).
Condition | Comment | |
Index | + | “It was mentally easier because there was less error with less fingers able to hit random buttons.” |
+ | “More of a freedom to look around and plan for the next key to tap rather than be constrained by my eyes.” | |
+ | “The other fingers couldn’t get in the way and I didn’t have to look at the space bar to press it.” | |
- | “The press detection feels a little unreliable.” | |
- | “I did not like the constraint of index fingers only because sometimes I would use my other fingers out of habit.” | |
- | “I still was looking down most of it because I don’t usually type with only my index fingers.” | |
IndexEye | + | “The eye-tracking is really neat, also definitely made me more accurate.” |
+ | “I was able to make the fewest mistakes in this condition.” | |
+ | “Only using index fingers made it easier to type the right keys.” | |
- | “Needing to look at the keys I want to type slows me down a little.” | |
- | “The lag from the eye-tracking.” | |
- | “It took me a bit to get used to not typing with all my fingers.” | |
Ten | + | “The speed of typing with unrestricted 10 fingers was fun.” |
+ | “A bit less visually cluttered.” | |
+ | “Being able to use my thumb to press space, and pinky to press shift.” | |
- | “I got frustrated a bit so I started using fewer fingers at times because it was easier.” | |
- | “I often found myself hitting space or another key when I didn’t want to.” | |
- | “Ring finger tracking inaccuracies.” | |
TenEye | + | “Definitely more accurate, was quicker typing because of that. Most false presses weren’t registered.” |
+ | “The flexibility of using all fingers.” | |
+ | “The eye-tracking made it harder to mis-type letters on the keyboard.” | |
- | “The HoloLens had trouble reading the ring and pinky fingers from most natural angles.” | |
- | “Was difficult to type quickly at first because I was getting used to the ‘look when you type’ system.” | |
- | “Not being able to use the space bar without looking at it.” |
4.5. Discussion
Our goals in Experiment 2 were to 1) examine if users could type faster in midair using all ten fingers, and 2) to see if eye-tracking could be used to reduce accidental key presses. However, we found users had a significantly higher entry rate in the Index and IndexEye conditions than in the Ten and TenEye conditions. Users also expressed an overwhelming preference for the index-finger conditions. Users performed poorly in the Ten condition as evidenced by lower entry rates than the two index-finger conditions and more backspaces per output character then all other conditions. Tactile feedback or improved tracking might have improved performance. Participants frequently complained in open feedback about the HoloLens’ hand-tracking. They also reported difficulty pressing keys using their ring fingers and frustration with a lack of responsiveness from the HoloLens.
Our hypothesis that users generally looked at the key they were attempting to press was supported by the the eye-tracking data gathered during the study. We examined this by noting how often people were looking near the key that they pressed in the Index and Ten conditions, which did not use eye-tracking to disable keys. If we had disabled keys far from their gaze point as in IndexEye and TenEye, key presses would have been excluded 32.9% and 36.0% of the time in the Index and Ten conditions respectively. This suggests that participants often looked where they were typing even when they were not required to do so. However, the IndexEye condition only excluded 10.4% of key presses. This notably lower amount may be simply because participants were more careful to match their gaze to the key they were pressing, lowering the number of unintentional key presses. However, in the TenEye condition, 36.7% of key presses occurred on disabled keys. This result is similar to the Ten condition and may show eye-tracking successfully filtered out unintentional key presses. However, it is important to note we do not know how many discarded key presses were due to unintentional presses being helpfully discarded versus key presses a participant wanted to make but that were discarded because they were not looking close enough to the key.
There is also other evidence that eye-tracking was successful. Firstly, responses indicated that eye-tracking generally did not increase eyestrain or physical fatigue experienced by users (Figure 13 bottom). Second, the addition of eye-tracking in TenEye significantly reduced the amount of backspaces per character (Figure 8 right). Lastly, character error rate was reduced in TenEye to levels similar to the index finger conditions. Overall, our results demonstrate that eye-tracking is an effective tool that should be considered when implementing ten-finger midair typing systems.
A common complaint during the eye-tracking conditions was needing to look at the space bar to trigger it. Since the space bar is large and thumbs often do not move far from it when typing with ten fingers, participants felt this unnecessarily slowed typing. In the future, allowing space bar presses regardless of gaze position may improve results. Another common complaint was that the keys were unresponsive or difficult to trigger. We are uncertain if this was due to hand-tracking failures or an unnecessarily deep threshold for triggering keys. Reducing the distance that participants needed to push keys would have made it easier for participants to trigger individual keys, potentially improving performance. However, a shorter distance threshold could also increase the frequency of accidental key presses. The half-second cooldown may have also limited text entry speed in some circumstance, but given all participants in our studies typed substantially slower than 24 WPM on average (i.e. two characters per second), we think this did not unduly slow typing. Additional work could investigate if a different cooldown could improve performance. However, similar to reducing the trigger distance, this could increase typing errors and actually slow typing.
Both experiments consisted of a single one-hour session. The majority of our participants had no prior mixed reality experience. For index-finger typing on a midair keyboard, it was reassuring to see that novices could achieve a functional level of text entry performance with little training. Furthermore, they did this on text requiring use of mixed case, punctuation, and numbers. However, it would be interesting to study virtual keyboard performance over a longer period of time to better understand whether better performance is attainable with practice. Long-term use would also allow exploring whether tuning a keyboard’s models or thresholds can improve performance for a specific user.
Our entry rates were undoubtedly limited by the deterministic nature of our keyboard. It is likely users could have typed substantially faster if they only had to aim approximately within a key or so of their desired letters. Powerful auto-correct might even be able to sift through the many false key triggers typical of ten-finger midair typing. However, even in the case a MR keyboard supporting fast typing via auto-correct, it would still need a fallback mode to support users entering exactly what they want. This can be necessary for certain types of content, or to avoid an “auto-correct trap” in which a user’s desired and correct text is repeatedly auto “corrected” into incorrect text. Our work here can help guide how an auto-correcting MR keyboard can best support fallback to deterministic input.
An interesting avenue for future research would be to use auxiliary sensors (e.g. the AR headset’s microphone or a smartwatch) to detect taps on a physical surface instead of, or in addition to, finger collision with the keys. This could improve tap detection accuracy of virtual keyboards co-located with physical surfaces. The tactile feedback of actual contact with a surface could mitigate the problem of unintentional key presses, improving ten-finger typing.
In our midair keyboard experiments, we used a user’s gaze location to make a hard decision that deactivated keys more than 2.6 cm away. This threshold was determined based on pilot testing and given the eye-tracking accuracy of our headset. It is possible a different threshold, or a more accurate eye-tracker, would improve performance of this feature. It would also be interesting to investigate improving location or tap detection accuracy via a softer, probabilistic fusion of the hand and eye-tracking signals. We have included in our supplementary materials555https://osf.io/q4n72 detailed participant log files from both experiments. This includes when participants triggered individual keys as well as continuously sampled observations from the hand and eye tracker for Experiment 2. This data could allow exploring potential gains from incorporating such changes into future midair keyboards. We have also included the source code for the keyboard applications we used in both experiments.
5. Conclusion
Experiment 1 showed that a midair keyboard had significantly higher entry rates and was preferred over keyboards co-located on a surface. Entry rates were 12.2, 9.1, and 7.8 WPM in the Midair, Table, and Wall conditions respectively. Questionnaire results also showed that our midair keyboard, which was arranged such that participants could rest their arms on a table, was preferred and not particularly tiring compared to the wall and table co-located keyboards. Participant feedback was generally positive about the tactile feedback and ergonomics of typing on a virtual keyboard located on a table. However, at present, the hand-tracking of a current commodity headset (HoloLens 2) probably need improvement to better support accurate deterministic typing near surfaces.
In Experiment 2, we further explored midair typing by implementing typing with ten fingers and using eye-tracking to help eliminate unintentional key presses and comparing it to index-finger typing. While we found eye-tracking reduced the need to correct errors, ten-finger typing in midair was still slower than typing with only the index fingers. Entry rates were 10.9, 9.4, 7.1, and 7.4 WPM in the Index, IndexEye, Ten, and TenEye conditions respectively. In situations where midair ten-finger typing is desired, our results suggest that the addition of eye-tracking can significantly reduce backspace use without negatively impacting either entry or final error rates. Questionnaire results also showed no significant negative impact of the addition of eye-tracking.
In summary, we found that using a commodity MR headset with no auxiliary devices or sensors, novices could type at a reasonable speed of around 11–12 WPM with a error rate below 1%. They did this on realistic real-world text involving mixed case, punctuation, and numbers. Furthermore, they did this on a deterministic keyboard without auto-correct. At present, of the options we tested, typing using just the index fingers on a floating midair keyboard appears to be the most preferable option.
Acknowledgements.
This material is based upon work supported by the NSF under Grant No. IIS-1909089. We thank Michigan Tech students in CS5641 from Spring 2024 for providing feedback on an initial draft of this paper.References
- (1)
- Adhikary and Vertanen (2021) Jiban Adhikary and Keith Vertanen. 2021. Text Entry in Virtual Environments using Speech and a Midair Keyboard. IEEE Transactions on Visualization and Computer Graphics 27, 5 (2021), 2648–2658. https://doi.org/10.1109/TVCG.2021.3067776
- Ahn et al. (2017) Sunggeun Ahn, Seongkook Heo, and Geehyuk Lee. 2017. Typing on a smartwatch for smart glasses. In Proceedings of the 2017 ACM International Conference on Interactive Surfaces and Spaces. 201–209.
- Baier and Burmester (2019) Monique Faye Baier and Michael Burmester. 2019. Not Just About the User: Acceptance of Speech Interaction in Public Spaces. In Proceedings of Mensch Und Computer 2019 (Hamburg, Germany) (MuC ’19). Association for Computing Machinery, New York, NY, USA, 349–359. https://doi.org/10.1145/3340764.3340801
- Boletsis and Kongsvik (2019) Costas Boletsis and Stian Kongsvik. 2019. Controller-based Text-input Techniques for Virtual Reality: An Empirical Comparison. International Journal of Virtual Reality 19, 3 (Oct. 2019), 2–15. https://doi.org/10.20870/IJVR.2019.19.3.2917
- Boring et al. (2009) Sebastian Boring, Marko Jurmu, and Andreas Butz. 2009. Scroll, tilt or move it: Using mobile phones to continuously control pointers on large public displays. In Proceedings of the 21st Annual Conference of the Australian Computer-Human Interaction Special Interest Group: Design: Open 24/7 (Melbourne, Australia) (OZCHI ’09). Association for Computing Machinery, New York, NY, USA, 161–168. https://doi.org/10.1145/1738826.1738853
- Bowman et al. (2002) Doug A. Bowman, Christopher J. Rhoton, and Marcio S. Pinho. 2002. Text Input Techniques for Immersive Virtual Environments: An Empirical Comparison. In Proceedings of the Human Factors and Ergonomics Society Annual Meeting, Vol. 46. 2154–2158. https://doi.org/10.1177/154193120204602611
- Cheng et al. (2022) Yi Fei Cheng, Tiffany Luong, Andreas Rene Fender, Paul Streli, and Christian Holz. 2022. ComforTable user interfaces: Surfaces reduce input error, time, and exertion for tabletop and mid-air user interfaces. In 2022 IEEE International Symposium on Mixed and Augmented Reality (ISMAR). IEEE, 150–159.
- Dudley et al. (2019) John Dudley, Hrvoje Benko, Daniel Wigdor, and Per Ola Kristensson. 2019. Performance Envelopes of Virtual Keyboard Text Input Strategies in Virtual Reality. In 2019 IEEE International Symposium on Mixed and Augmented Reality (ISMAR). 289–300. https://doi.org/10.1109/ISMAR.2019.00027
- Dudley et al. (2018) John J. Dudley, Keith Vertanen, and Per Ola Kristensson. 2018. Fast and Precise Touch-Based Text Entry for Head-Mounted Augmented Reality with Variable Occlusion. ACM Transactions on Computer-Human Interaction (TOCHI) 25, 6, Article 30 (12 2018), 40 pages. https://doi.org/10.1145/3232163
- Feit et al. (2015) Anna Maria Feit, Srinath Sridhar, Christian Theobalt, and Antti Oulasvirta. 2015. Investigating multi-finger gestures for mid-air text entry. ACM womENcourage (2015).
- Feng et al. (2021) Wenxin Feng, Jiangnan Zou, Andrew Kurauchi, Carlos H Morimoto, and Margrit Betke. 2021. HGaze Typing: Head-Gesture Assisted Gaze Typing. In ACM Symposium on Eye Tracking Research and Applications (Virtual Event, Germany) (ETRA ’21 Full Papers). Association for Computing Machinery, New York, NY, USA, Article 11, 11 pages. https://doi.org/10.1145/3448017.3457379
- Findlater et al. (2011) Leah Findlater, Jacob O. Wobbrock, and Daniel Wigdor. 2011. Typing on flat glass: examining ten-finger expert typing patterns on touch surfaces. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Vancouver, BC, Canada) (CHI ’11). Association for Computing Machinery, New York, NY, USA, 2453–2462. https://doi.org/10.1145/1978942.1979301
- Foy et al. (2021) Conor R. Foy, John J. Dudley, Aakar Gupta, Hrvoje Benko, and Per Ola Kristensson. 2021. Understanding, Detecting and Mitigating the Effects of Coactivations in Ten-Finger Mid-Air Typing in Virtual Reality. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (, Yokohama, Japan,) (CHI ’21). Association for Computing Machinery, New York, NY, USA, Article 287, 11 pages. https://doi.org/10.1145/3411764.3445671
- Gizatdinova et al. (2018) Yulia Gizatdinova, Oleg Špakov, Outi Tuisku, Matthew Turk, and Veikko Surakka. 2018. Gaze and head pointing for hands-free text entry: applicability to ultra-small virtual keyboards. In Proceedings of the 2018 ACM Symposium on Eye Tracking Research & Applications (Warsaw, Poland) (ETRA ’18). Association for Computing Machinery, New York, NY, USA, Article 14, 9 pages. https://doi.org/10.1145/3204493.3204539
- Grubert et al. (2018) Jens Grubert, Lukas Witzani, Eyal Ofek, Michel Pahud, Matthias Kranz, and Per Ola Kristensson. 2018. Text Entry in Immersive Head-Mounted Display-Based Virtual Reality Using Standard Keyboards. In 2018 IEEE Conference on Virtual Reality and 3D User Interfaces (VR). 159–166. https://doi.org/10.1109/VR.2018.8446059
- Gu et al. (2016) Xiaochi Gu, Yifei Zhang, Weize Sun, Yuanzhe Bian, Dao Zhou, and Per Ola Kristensson. 2016. Dexmo: An inexpensive and lightweight mechanical exoskeleton for motion capture and force feedback in VR. In Proceedings of the 2016 CHI Conference on Human Factors in Computing Systems. 1991–1995.
- Gupta et al. (2020) Aakar Gupta, Majed Samad, Kenrick Kin, Per Ola Kristensson, and Hrvoje Benko. 2020. Investigating Remote Tactile Feedback for Mid-Air Text-Entry in Virtual Reality. In 2020 IEEE International Symposium on Mixed and Augmented Reality (ISMAR). 350–360. https://doi.org/10.1109/ISMAR50242.2020.00062
- He et al. (2022) Zhenyi He, Christof Lutteroth, and Ken Perlin. 2022. TapGazer: Text Entry with Finger Tapping and Gaze-directed Word Selection. In Proceedings of Conference on Human Factors in Computing Systems. Article 337, 16 pages. https://doi.org/10.1145/3491102.3501838
- Hincapié-Ramos et al. (2014) Juan David Hincapié-Ramos, Xiang Guo, Paymahn Moghadasian, and Pourang Irani. 2014. Consumed endurance: a metric to quantify arm fatigue of mid-air interactions. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Toronto, Ontario, Canada) (CHI ’14). Association for Computing Machinery, New York, NY, USA, 1063–1072. https://doi.org/10.1145/2556288.2557130
- Hu et al. (2024) Jinghui Hu, John J. Dudley, and Per Ola Kristensson. 2024. SkiMR: Dwell-free Eye Typing in Mixed Reality. In 2024 IEEE Conference Virtual Reality and 3D User Interfaces (VR). 439–449. https://doi.org/10.1109/VR58804.2024.00065
- Kim et al. (2023) Taejun Kim, Amy Karlson, Aakar Gupta, Tovi Grossman, Jason Wu, Parastoo Abtahi, Christopher Collins, Michael Glueck, and Hemant Bhaskar Surale. 2023. STAR: Smartphone-analogous Typing in Augmented Reality. In Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology (San Francisco, CA, USA) (UIST ’23). Association for Computing Machinery, New York, NY, USA, Article 116, 13 pages. https://doi.org/10.1145/3586183.3606803
- Knierim et al. (2018) Pascal Knierim, Valentin Schwind, Anna Maria Feit, Florian Nieuwenhuizen, and Niels Henze. 2018. Physical Keyboards in Virtual Reality: Analysis of Typing Performance and Effects of Avatar Hands. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems (Montreal, QC, Canada) (CHI ’18). Association for Computing Machinery, New York, NY, USA, Article 345, 9 pages. https://doi.org/10.1145/3173574.3173919
- Kristensson and Vertanen (2012) Per Ola Kristensson and Keith Vertanen. 2012. The Potential of Dwell-Free Eye-Typing for Fast Assistive Gaze Communication. In ETRA ’12: Proceedings of the ACM Symposium on Eye-Tracking Research and Applications (Santa Barbara, CA). 241–244. https://doi.org/10.1145/2168556.2168605
- Kristensson and Zhai (2004) Per-Ola Kristensson and Shumin Zhai. 2004. SHARK2: A Large Vocabulary Shorthand Writing System for Pen-based Computers. In Proceedings of the 17th annual ACM symposium on User interface software and technology (Santa Fe, NM, USA) (UIST ’04). ACM, ACM, New York, NY, USA, 43–52. https://doi.org/10.1145/1029632.1029640
- Lee and Woo (2003) Minkyung Lee and Woontack Woo. 2003. ARKB: 3D vision-based Augmented Reality Keyboard. In Proceeding of the 13th International Conference on Artificial Reality and Telexistence (Tokyo, Japan).
- Lu et al. (2021) Xueshi Lu, Difeng Yu, Hai-Ning Liang, and Jorge Goncalves. 2021. iText: Hands-free Text Entry on an Imaginary Keyboard for Augmented Reality Systems. In Proceedings of the 34th Annual ACM Symposium on User Interface Software and Technology (Virtual Event, USA) (UIST ’21). Association for Computing Machinery, New York, NY, USA, 815–825. https://doi.org/10.1145/3472749.3474788
- Lystbæk et al. (2022) Mathias N. Lystbæk, Ken Pfeuffer, Jens Emil Sloth Grønbæk, and Hans Gellersen. 2022. Exploring Gaze for Assisting Freehand Selection-based Text Entry in AR. In Proceedings of the ACM on Human-Computer Interaction, Vol. 6. Article 141, 16 pages. https://doi.org/10.1145/3530882
- Malik et al. (2005) Shahzad Malik, Abhishek Ranjan, and Ravin Balakrishnan. 2005. Interacting with large displays from a distance with vision-tracked multi-finger gestural input. In Proceedings of the 18th Annual ACM Symposium on User Interface Software and Technology (Seattle, WA, USA) (UIST ’05). Association for Computing Machinery, New York, NY, USA, 43–52. https://doi.org/10.1145/1095034.1095042
- Markussen et al. (2013) Anders Markussen, Mikkel R Jakobsen, and Kasper Hornbæk. 2013. Selection-based Mid-air Text Entry on Large Displays. In IFIP Conference on Human-Computer Interaction – INTERACT 2013. Springer, Berlin, Heidelberg, 401–418. https://doi.org/10.1007/978-3-642-40483-2_28
- Markussen et al. (2014) Anders Markussen, Mikkel Rønne Jakobsen, and Kasper Hornbæk. 2014. Vulture: A Mid-Air Word-Gesture Keyboard. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Toronto, Ontario, Canada) (CHI ’14). ACM, New York, NY, USA, 1073–1082. https://doi.org/10.1145/2556288.2556964
- McCaul and Sutherland (2004) Barry McCaul and Alistair Sutherland. 2004. Predictive text entry in immersive environments. In IEEE Virtual Reality 2004. IEEE, 241–242.
- McGill et al. (2015) Mark McGill, Daniel Boland, Roderick Murray-Smith, and Stephen Brewster. 2015. A Dose of Reality: Overcoming Usability Challenges in VR Head-Mounted Displays. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Seoul, Republic of Korea) (CHI ’15). ACM, New York, NY, USA, 2143–2152. https://doi.org/10.1145/2702123.2702382
- Pezent et al. (2019) Evan Pezent, Ali Israr, Majed Samad, Shea Robinson, Priyanshu Agarwal, Hrvoje Benko, and Nick Colonnese. 2019. Tasbi: Multisensory Squeeze and Vibrotactile Wrist Haptics for Augmented and Virtual Reality. In 2019 IEEE World Haptics Conference (WHC). 1–6. https://doi.org/10.1109/WHC.2019.8816098
- Pfeuffer et al. (2020) Ken Pfeuffer, Lukas Mecke, Sarah Delgado Rodriguez, Mariam Hassib, Hannah Maier, and Florian Alt. 2020. Empirical Evaluation of Gaze-enhanced Menus in Virtual Reality. In Proceedings of the 26th ACM Symposium on Virtual Reality Software and Technology (Virtual Event, Canada) (VRST ’20). Association for Computing Machinery, New York, NY, USA, Article 20, 11 pages. https://doi.org/10.1145/3385956.3418962
- Pick et al. (2016) Sebastian Pick, Andrew S. Puika, and Torsten W. Kuhlen. 2016. SWIFTER: Design and evaluation of a speech-based text input metaphor for immersive virtual environments. In 2016 IEEE Symposium on 3D User Interfaces (3DUI). 109–112. https://doi.org/10.1109/3DUI.2016.7460039
- Rabin and Gordon (2004) Ely Rabin and Andrew Gordon. 2004. Tactile feedback contributes to consistency of finger movements during typing. Experimental Brain Research 155 (2004), 362–369. https://doi.org/10.1007/s00221-003-1736-6
- Richardson et al. (2024) Mark Richardson, Fadi Botros, Yangyang Shi, Pinhao Guo, Bradford J Snow, Linguang Zhang, Jingming Dong, Keith Vertanen, Shugao Ma, and Robert Wang. 2024. StegoType: Surface Typing from Egocentric Cameras. In Proceedings of the 37th Annual ACM Symposium on User Interface Software and Technology. 1–14.
- Richardson et al. (2020) Mark Richardson, Matt Durasoff, and Robert Wang. 2020. Decoding surface touch typing from hand-tracking. In Proceedings of the 33rd annual ACM symposium on user interface software and technology. 686–696.
- Roeber et al. (2003) Helena Roeber, John Bacus, and Carlo Tomasi. 2003. Typing in thin air: the canesta projection keyboard - a new method of interaction with electronic devices. In CHI ’03 Extended Abstracts on Human Factors in Computing Systems (Ft. Lauderdale, Florida, USA) (CHI EA ’03). Association for Computing Machinery, New York, NY, USA, 712–713. https://doi.org/10.1145/765891.765944
- Sakashita et al. (2019) Mose Sakashita, Satoshi Hashizume, and Yoichi Ochiai. 2019. Wrist-mounted haptic feedback for support of virtual reality in combination with electrical muscle stimulation and hanger reflex. In Human-Computer Interaction. Recognition and Interaction Technologies: Thematic Area, HCI 2019, Held as Part of the 21st HCI International Conference, HCII 2019, Orlando, FL, USA, July 26–31, 2019, Proceedings, Part II 21. Springer, 544–553.
- Speicher et al. (2018a) Marco Speicher, Anna Maria Feit, Pascal Ziegler, and Antonio Krüger. 2018a. Selection-Based Text Entry in Virtual Reality. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems (Montreal QC, Canada) (CHI ’18). ACM, New York, NY, USA, 1–13. https://doi.org/10.1145/3173574.3174221
- Speicher et al. (2018b) Marco Speicher, Anna Maria Feit, Pascal Ziegler, and Antonio Krüger. 2018b. Selection-based Text Entry in Virtual Reality. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems (Montreal, QC, Canada) (CHI ’18). Association for Computing Machinery, New York, NY, USA, Article 647, 13 pages. https://doi.org/10.1145/3173574.3174221
- Sridhar et al. (2015) Srinath Sridhar, Anna Maria Feit, Christian Theobalt, and Antti Oulasvirta. 2015. Investigating the Dexterity of Multi-Finger Input for Mid-Air Text Entry. In Proceedings of the 33rd Annual ACM Conference on Human Factors in Computing Systems (Seoul, Republic of Korea) (CHI ’15). Association for Computing Machinery, New York, NY, USA, 3643–3652. https://doi.org/10.1145/2702123.2702136
- Vertanen and Kristensson (2011) Keith Vertanen and Per Ola Kristensson. 2011. A versatile dataset for text entry evaluations based on genuine mobile emails. In Proceedings of the 13th International Conference on Human Computer Interaction with Mobile Devices and Services (Stockholm, Sweden) (MobileHCI ’11). Association for Computing Machinery, New York, NY, USA, 295–298. https://doi.org/10.1145/2037373.2037418
- Walker et al. (2017) James Walker, Bochao Li, Keith Vertanen, and Scott Kuhl. 2017. Efficient Typing on a Visually Occluded Physical Keyboard. In Proceedings of the 2017 CHI Conference on Human Factors in Computing Systems (Denver, Colorado, USA) (CHI ’17). Association for Computing Machinery, New York, NY, USA, 5457–5461. https://doi.org/10.1145/3025453.3025783
- Wobbrock (2007) Jacob O Wobbrock. 2007. Measures of text entry performance. Text entry systems: Mobility, accessibility, universality (2007), 47–74.
- Xu et al. (2019) W. Xu, H. Liang, A. He, and Z. Wang. 2019. Pointing and Selection Methods for Text Entry in Augmented Reality Head Mounted Displays. In 2019 IEEE International Symposium on Mixed and Augmented Reality (ISMAR). 279–288. https://doi.org/10.1109/ISMAR.2019.00026
- Xu et al. (2019a) Wenge Xu, Hai-Ning Liang, Anqi He, and Zifan Wang. 2019a. Pointing and selection methods for text entry in augmented reality head mounted displays. In 2019 IEEE International Symposium on Mixed and Augmented Reality (ISMAR). IEEE, 279–288.
- Xu et al. (2019b) Wenge Xu, Hai-Ning Liang, Yuxuan Zhao, Tianyu Zhang, Difeng Yu, and Diego Monteiro. 2019b. RingText: Dwell-free and hands-free Text Entry for Mobile Head-Mounted Displays using Head Motions. IEEE transactions on visualization and computer graphics 25, 5 (2019), 1991–2001.
- Yi et al. (2015) Xin Yi, Chun Yu, Mingrui Zhang, Sida Gao, Ke Sun, and Yuanchun Shi. 2015. ATK: Enabling Ten-Finger Freehand Typing in Air Based on 3D Hand Tracking Data. In Proceedings of the 28th Annual ACM Symposium on User Interface Software and Technology (Charlotte, NC, USA) (UIST ’15). ACM, New York, NY, USA, 539–548. https://doi.org/10.1145/2807442.2807504
- Yu et al. (2017) Chun Yu, Yizheng Gu, Zhican Tang, Xin Yi, Hengliang Luo, and Yuanchun Shi. 2017. Tap, dwell or gesture?: Exploring head-based text entry techniques for HMDs. In Proceedings of Conference on Human Factors in Computing Systems. 4479–2288. https://doi.org/10.1145/3025453.3025964
- Zhai and Kristensson (2012) Shumin Zhai and Per Ola Kristensson. 2012. The Word-gesture Keyboard: Reimagining Keyboard Interaction. Communications of the ACM 55, 9 (month 2012), 91–101. https://doi.org/10.1145/2330667.2330689

### v3

@@
-The Impact of Surface Co-location and Eye-tracking on Mixed Reality Typing
+Exploring Keyboard Positioning and Ten-Finger Typing in Mixed Reality
 Abstract.
 Accuracy and speed are pivotal when typing.
-We hypothesized that the lack of tactile feedback on midair mixed reality keyboards may adversely impact performance.
-Our first experiment investigated providing tactile feedback to users typing in mixed reality by co-locating the virtual keyboard on a table or a wall.
-The keyboard was deterministic (without auto-correct), allowed mixed case typing with symbols, and relied only on the hand-tracking provided by a commodity headset’s egocentric cameras.
+Mixed reality typing is typically performed by typing on a midair keyboard with your index fingers.
+This deprives users of both the tactile feedback available on physical devices and the ability to press keys with the most convenient finger.
+Our first experiment investigated providing tactile feedback by positioning the virtual keyboard on a table or wall.
+The keyboard was deterministic (without auto-correct), supported mixed case typing with symbols, and relied only on the hand-tracking provided by a commodity headset’s egocentric cameras.
 Users preferred and had the highest entry rate of 12 words-per-minute using a midair keyboard.
 Error rates were similar in all conditions.
-Based on user feedback, our second experiment explored ten-finger typing.
-We used a novel eye-tracking technique to mitigate accidental key presses.
-This technique was preferred for ten-finger typing and halved the number of times backspace was pressed.
-However, participants were faster using only their index fingers without eye-tracking at 11 words-per-minute.
+Our second experiment explored ten-finger typing and used a novel eye-tracking technique to avoid accidental key presses.
+This technique was preferred for ten-finger typing and halved corrections.
+However, participants were faster using their index fingers without eye-tracking at 11 words-per-minute.
 1. Introduction
 Mixed reality (MR) offers the ability to project virtual objects onto the real world and interact with them.
-As MR technology develops, it may begin to displace other portable computing devices such as cell phones and laptops.
-Users might wish to utilize MR to perform everyday tasks.
+As MR technology develops, users might wish to utilize MR to perform everyday tasks.
 Typing is critical to efficiently performing common tasks such as writing emails, browsing the web, and creating documents.
 Unlike a traditional keyboard where you can feel the locations of the keys and when they are actuated, MR systems often employ midair keyboards which lack tactile feedback.
-The loss of tactile feedback can impair users’ accuracy when typing (Rabin and Gordon, 2004), potentially reducing the utility of MR compared to conventional devices for supporting common, everyday computing activities.
-Speech input offers an alternative to typing but may not always be socially acceptable (Baier and Burmester, 2019), and recognition of difficult words can be a problem (Adhikary and Vertanen, 2021).
+The loss of tactile feedback has been shown to impair users’ accuracy when typing (Rabin and Gordon, 2004).
+Further, MR systems typically restrict typing to just a user’s index fingers.
+These issues can potentially reduce the utility of MR compared to conventional devices for supporting common, everyday computing activities.
+Speech input offers an alternative to typing but may not always be socially acceptable (Baier and Burmester, 2019), and recognition of difficult words can be a problem (Adhikary and Vertanen, 2021a).
 Auxiliary devices such as physical keyboards (McGill et al., 2015; Walker et al., 2017; Knierim et al., 2018; Grubert et al., 2018), controllers (Speicher et al., 2018a; Xu et al., 2019a), gloves (Markussen et al., 2014; McCaul and Sutherland, 2004; He et al., 2022), or watches (Ahn et al., 2017), could help but require carrying additional devices.
@@
 However, such methods introduce a learning curve that may be a barrier to MR adoption.
-Compared to past work, we focus on enabling mixed reality typing without: 1) requiring users to learn a new interaction technique, 2) requiring devices or sensors aside from the headset, and 3) using a language model to compensate for noisy input.
-We conjectured providing tactile feedback would improve typing in MR.
-In Experiment 1, we aligned a virtual keyboard with a wall or table.
-This adds tactile feedback to keypresses and could be used whenever a flat surface is present.
-Although surface co-location may not be feasible in all situations, people are frequently in human-made environments where flat surfaces are prevalent.
-We compared the surface aligned keyboards with one floating in midair.
+Such methods also usually rely on a language model which restricts input to predictable natural language text.
+Compared to past work, we focus on mixed reality typing that: 1) leverages users’ existing knowledge of tapping on an virtual keyboard, 2) uses only the sensors built into the headset, and 3) allows input of arbitrary character sequences (i.e.
+without relying on a language model to compensate for noisy input).
+Tactile feedback of a virtual keyboard on a table has been shown to improve performance (Dudley et al., 2019).
+What hasn’t been shown is whether this is feasible using only the tracking available from the egocentric sensors on the headset and without relying on auto-correct.
+Further, only a keyboard positioned on a horizontal surface has been explored previously.
+In Experiment 1, we explored three keyboard positions: midair, table aligned, and wall aligned.
+The table and wall surfaces add tactile feedback to keypresses and could be used whenever a flat surface is present.
+Although locating a keyboard on a surface may not be feasible in all situations, people are frequently in human-made environments where flat surfaces are prevalent.
 Participants typed using only their index fingers.
+They typed faster in midair at 12.2 words-per-minute (WPM) compared to on a table (9.1 WPM) or a wall (8.8 WPM).
 We found participants preferred the midair keyboard compared to keyboards near a table or wall.
-They typed faster in midair at 12.2 words-per-minute (WPM) compared to on a table (9.1 WPM) or a wall (8.8 WPM).
 Experiment 2 focused on ten-finger typing, a feature requested by many participants in our first study.
@@
 mixed case with punctuation and numbers) accurately and at a reasonable speed of 11–12 words-per-minute on a deterministic midair keyboard.
-User feedback was in general positive for the tactile feedback and ergonomics offered by virtual keyboards located near a surface.
-However near surfaces, accurate hand-tracking was a challenge for a current commercial headset.
+While many users liked the tactile feedback afforded by a keyboard on a table or wall, we show midair typing provides higher performance on a current commercial headset due to challenges of hand-tracking near a surface.
 Our novel eye-tracking based filter feature reduced unintentional key presses in midair ten-finger typing.
-We release detailed logs from our studies and the source code of our HoloLens keyboard experimental applications.
+We release detailed logs from our studies and the source code of our HoloLens111HoloLens is an augmented reality headset released by Microsoft in 2016.
+A second version of the headset (and the one we used) was released in 2019.
+experimental applications.
 2. Related Work
-If mixed reality is going to be used to perform everyday computing tasks, text input needs to be efficient.
+2.1. Midair virtual keyboards
+For mixed reality to be useful for everyday computing tasks, text input needs to be efficient.
 Many solutions have been proposed to address this need.
-A midair virtual keyboard that allows users to type directly on it with their hands is both familiar and easy-to-learn.
-Early systems such as ARKB (Lee and Woo, 2003) relied on external cameras to track a user’s hands over a virtual keyboard.
+A midair virtual keyboard allowing users to type directly with their hands is both familiar and easy-to-learn.
+Early systems such as ARKB (Lee and Woo, 2003) relied on external cameras to track a user’s hands over the keyboard.
+Similar to our first study, ARKB used printed fiducial markers to position a keyboard horizontally and slightly above a table.
+No user study results were reported.
 Markussen et al.
-(Markussen et al., 2013) used an array of 24 external cameras to track a user’s hands.
-Users directly selected keys on a projected QWERTY keyboard, or used ambiguous gestures to specify keys.
+(2013) also relied on external cameras for hand-tracking.
+Users either directly selected keys or used ambiguous gestures to specify keys.
+The keyboard was projected vertically in front of users on a large display.
 Direct typing was the fastest at 13 WPM after six sessions.
 Dudley et al.
-(Dudley et al., 2018) investigated midair virtual keyboard typing on a HoloLens 1.
+(2018) investigated midair typing on a HoloLens 1.
+The keyboard was located in midair slightly below a user’s eyes and was tilted 20 degrees away from vertical.
 Similar to our work, they relied only on the built-in headset sensors.
 Despite the HoloLens 1 only tracking a user’s wrist position, users typed at 18 WPM on an auto-correcting keyboard with word predictions.
-In comparison, we used a HoloLens 2 with full hand-tracking and focused on input of more difficult mixed case text with symbols and numbers.
-Cheng et al.
-(Cheng et al., 2022) explored if sitting at a table improved users’ ability to select and drag targets in virtual reality (VR).
-Users’ arms and finger were tracked via external cameras.
-They found the presence of a table improved user accuracy and reduced fatigue of horizontal and vertical midair interfaces.
-In this work, we also chose to use a table as it is a common everyday object.
-However, we focus on MR and the more accuracy demanding task of keyboard typing.
+In comparison to past work, we used a HoloLens 2 with full hand-tracking, eschewed the use of external cameras, and focused on the input of more difficult mixed case text with symbols and numbers.
+2.2. Physical keyboards in VR and AR
+Various work has explored typing on a physical keyboard while wearing a VR or AR headset.
+McGill et al.
+(2015) had users type: 1) without a headset, 2) wearing a headset with no view of the keyboard or a user’s hands, 3) wearing a headset with a view of the keyboard and a user’s hands blended in, or 4) wearing a headset with a view of the real-world.
+Users were fastest without the headset (59 WPM) and slowest when they had no view of the keyboard or their hands (24 WPM).
+Blending or full real-world visual feedback provided similar performance at 39 and 37 WPM respectively.
+Rather than blending real-world video into VR, Walker et al.
+(2017) highlighted keys on a virtual keyboard as a user typed.
+Coupled with auto-correct, users typed in VR with only a virtual keyboard at 44 WPM.
+Chiossi et al.
+(2024) had users type in AR, VR, and augmented virtuality (where real-world imagery is displayed in VR (McGill et al., 2015)).
+Users typed at similar speeds of 34.9, 32.5, and 34.5 WPM respectively.
+While typing on a physical keyboard can be fast and support rich text input, users may not want to carry an extra device solely for text entry.
+As such, we focus on using a virtual keyboard.
+2.3. Tactile feedback
 While keyboard typing is a familiar interaction, the lack of tactile feedback on a midair keyboard can be a challenge.
@@
 Dudley et al.
-(Dudley et al., 2019) had users type with two or ten fingers on virtual keyboards located near a surface or in midair.
-Users’ hands were tracked via an array of external cameras and markers on the hands.
-The system used an auto-correct algorithm based on the system knowing the target text.
-Users typed significantly faster on the surface keyboards with either two fingers (56 WPM versus 42 WPM) or all ten fingers (52 WPM versus 35 WPM).
+(2019) had users type with two or ten fingers on virtual keyboards located near a surface or in midair.
+Users’ hands were tracked via an external camera array and markers on their hands.
+Users were seated at a desk and the midair keyboard was tilted slightly above horizontal.
+The system used an auto-correct algorithm based on the known target text.
+Users typed faster on the surface keyboards with either two fingers (56 WPM versus 42 WPM) or all ten fingers (52 WPM versus 35 WPM).
+Our work focuses on input without auto-correct and without knowing the target text.
 Richardson et al.
-(Richardson et al., 2020) also investigated surface typing by tracking users’ hands via external cameras.
+(2020) also investigated surface typing by tracking users’ hands via external cameras.
 Users typed phrases quickly at 73 WPM.
-Using a neural network hand motion model and a language model, they found recorded typing sequences could be decoded at a low error rate of 2%.
-In recent followup work, Richardson et al.
-(Richardson et al., 2024) used only the egocentric cameras of a VR headset to infer keypresses on a virtual keyboard situated on a desk.
+Using a neural network hand motion model and a language model, they typed sequences could be decoded at a low error rate of 2%.
+In followup work, Richardson et al.
+(2024) used only the egocentric cameras of a VR headset to infer keypresses on a virtual keyboard situated on a desk.
 Users typed at 44 WPM with an uncorrected error rate of 7%.
@@
 Indeed, on non-language text, offline experiments showed a much high error rate of 16% compared to 5% on English.
-Similar to (Richardson et al., 2024), we utilize the hand-tracking available from the egocentric cameras of a commodity headset.
+Similar to Richardson et al.
+(2024), we utilize the hand-tracking available from the egocentric cameras of a commodity headset.
 However, we rely solely on a human-in-the-loop process where the user observes their inferred finger locations and uses this feedback to supervise their typing.
@@
 typing a password).
+2.4. Interaction context
+Mixed reality offers the possibly of interaction in many real-world contexts, for example while sitting, standing, or walking.
+Cheng et al.
+(2022) explored if sitting at a table improved users’ ability to select and drag targets in VR.
+Users’ arms and finger were tracked via external cameras.
+They found having a table improved user accuracy and reduced fatigue of horizontal and vertical midair interfaces.
+Dudley et al.
+(2019) controlled for posture by having users sit at a desk while they typed on a virtual keyboard.
+We also chose to investigate seated interaction as this is a common context in which people do productivity related text entry.
+Further, given our chosen task of entering rich text without auto-correct, we wanted to reduce physical fatigue and user motion that could make hand-tracking interaction more challenging.
+2.5. Alternatives to QWERTY typing
 Moving somewhat away from the metaphor of tapping on a keyboard, a word-gesture keyboard (Kristensson and Zhai, 2004; Zhai and Kristensson, 2012) allows writing by swiping through a word’s letters.
@@
 Early work allowed users wearing a glove to write words in VR by selecting different sets of letters by flexing their finger (McCaul and Sutherland, 2004).
-A similar recent system is TapGazer (He et al., 2022) which uses a glove to detect which finger a user is tapping with disambiguation via eye gaze or additional taps.
+A similar recent system is TapGazer (He et al., 2022) which uses a glove to detect finger taps with disambiguation via eye gaze or additional taps.
 Other approaches for text input in VR and AR have explored input methods other than hand gestures.
@@
 While Gizatdinova et al.
-(Gizatdinova et al., 2018) found head-based typing was effective on small keyboards, users thought it felt unnatural and they preferred gaze-based typing.
+(2018) found head-based typing was effective on small keyboards, users thought it felt unnatural and they preferred gaze-based typing.
 Gaze-based typing in VR has been investigated both as the sole input method (Lu et al., 2021) and in combination with head gestures (Feng et al., 2021).
@@
 Using this approach, Hu et al.
-(Hu et al., 2024) found users wrote at 12 WPM using a HoloLens 2 headset.
+(2024) found users wrote at 12 WPM using a HoloLens 2 headset.
 While potentially fast, dwell-free typing requires user learning and may not work for hard-to-predict text.
+2.6. Speech input
 Using speech input can help reduce MR challenges such as robustly sensing users hands and accurately displaying visual feedback.
 In early work, Bowman et al.
-(Bowman et al., 2002) explored VR text input using Wizard of Oz speech recognition and input via a glove, keyboard, or tablet.
+(2002) explored VR text input using Wizard of Oz speech recognition and input via a glove, keyboard, or tablet.
 Speech recognition was fastest at around 13 WPM.
 In SWIFTER (Pick et al., 2016), users wrote in a virtual environment at 24 WPM using speech with correction via a hand controller.
-Similar to our system, Adhikary and Vertanen (Adhikary and Vertanen, 2021) eschewed the use of auxiliary devices and relied solely on speech and hand-tracking for VR text input.
+Similar to our system, Adhikary and Vertanen (2021a) avoided auxiliary devices and relied solely on speech and hand-tracking for VR text input.
 Speaking and then correcting via hand gestures was the fastest at 28 WPM.
@@
 usernames and passwords).
-Most existing MR typing solutions, while often allowing typing with both hands, typically only support typing with the index fingers (Lystbæk et al., 2022; Dudley et al., 2018; Adhikary and Vertanen, 2021).
-Systems that do allow use of all fingers generally require an auxiliary input device (He et al., 2022), external sensing device (Yi et al., 2015), or physical keyboard (McGill et al., 2015; Walker et al., 2017; Knierim et al., 2018; Grubert et al., 2018).
-Ten-finger typing is challenging as the dexterity limitations of the human hand can make some multi-finger gestures difficult to perform without the tactile assistance of a physical keyboard (Sridhar et al., 2015).
-Additionally, ten-finger typing on virtual keyboards can lead to accidental inputs and key coactivations (Foy et al., 2021).
+2.7. Ten finger interaction
+While existing MR systems often allow typing with both hands, typically they only support use of the index fingers (Lystbæk et al., 2022; Dudley et al., 2018; Adhikary and Vertanen, 2021a).
+Systems that allow use of all fingers generally require an auxiliary input device (He et al., 2022), external sensing device (Yi et al., 2015), or physical keyboard (McGill et al., 2015; Walker et al., 2017; Knierim et al., 2018; Grubert et al., 2018).
+Ten-finger typing is challenging as the dexterity limitations of the human hand can make some multi-finger gestures difficult without the tactile assistance of a physical keyboard (Sridhar et al., 2015).
+Additionally, ten-finger typing on virtual keyboards can lead to accidental key coactivations (Foy et al., 2021).
 While multi-finger gestures have been used for navigating MR interfaces, the small movements required for typing may be difficult for vision-based hand-tracking systems to recognize (Malik et al., 2005).
-To try and improve ten-finger typing, in Experiment 2, we use eye-tracking to filter out key presses not located near a user’s gaze.
-Past work shows that users will naturally tend to look at the keys they are pressing (Lystbæk et al., 2022).
+Past work has showed that users naturally look at the keys they are pressing (Lystbæk et al., 2022).
+Thus in Experiment 2, we explore using eye-tracking to filter out key presses not near a user’s gaze.
 3. Experiment 1
 In our first experiment, participants wore a HoloLens 2 headset and typed sentences on a virtual keyboard using their index fingers.
-We measured user performance and preference for typing on a midair keyboard and keyboards co-located with a table and with a wall.
-Our primary goal was to assess the impact of the tactile feedback offered by the surface co-location.
+We measured user performance and preference for typing on a midair keyboard and keyboards located on a table and on a wall.
+Our primary goal was to assess the impact of different keyboard positions, including placement on a horizontal and vertical surface.
 We opted not to include a physical keyboard condition since typing on physical keyboards has been well-studied and our focus is on on-the-go applications where all sensing is performed by the headset.
@@
 The experiment consisted of three conditions (depicted in Figure 1):
-Table — The keyboard was placed just above a table, using a QR code on the table to ensure consistent spatial placement.
-The keyboard was placed horizontally, such that the back of the keyboard was parallel with the table.
-Wall — The keyboard was placed slightly in front of a wall, using the same QR code as a spatial anchor.
-The keyboard was placed vertically, such that the back of the keyboard was parallel with the wall.
-We hypothesized the tactile feedback conditions (i.e.
-Wall and Table) would perform better than the non-tactile feedback condition (i.e.
+Midair — The keyboard was placed in midair in front of the participant.
+The keyboard was horizontally centered above the table with the bottom of the spacebar key being 19.0 cm from the front of the table and 4.5 cm above the table.
+This position made it possible for users to rest their forearms against the edge of the table while typing if they so desired.
+The keyboard was angled 15 degrees above horizontal.
+In pilot testing, we found this tilt made it more comfortable to type while seated compared to a vertical or completely horizontal keyboard.
+A similar tilt was used in a study where participants typed on or above a surface while seated (Dudley et al., 2019) (c.f.
+Figure 5).
+Table — The keyboard was placed just above a table.
+The keyboard was horizontally centered on the table with the bottom of the spacebar key being 28.5 cm from the front of the table.
+This position made it possible for users to rest their arms on the table as they typed if they so desired.
+Wall — The keyboard was placed slightly in front of a wall.
+The keyboard was horizontally centered on the table with the bottom of the spacebar being 16.0 cm above the table.
+This position coupled with the small size of the table made it possible for users to rest their arms on the table as they typed if they so desired.
+We hypothesized the keyboards positions on surfaces (i.e.
+Wall and Table) would perform better than the baseline condition (i.e.
 Midair).
@@
 This allowed us to investigate whether this impacted hand-tracking accuracy.
-In all conditions, participants used their index fingers on one or both hands to enter text.
+In all conditions, participants typed with their index fingers on one or both hands.
 This choice prioritizes ecological validity by allowing participants to type as they would in the real world.
-Participants also could choose to rest their arms on a table as needed to potentially reduce fatigue.
-Our keyboard was deterministic and did not auto-correct or predict words.
+Our keyboard was deterministic and did not perform auto-correction or predict words.
 We did this to understand the raw performance of typing as might be required for text not easily predicted by a language model.
@@
 A key was triggered when either index finger entered the key’s bounding area, pushed the key’s face a certain depth, and exited the same key’s bounding area.
-On a key press, we played a button sound effect and displayed a white border around the key.
+On a key press, we played a sound effect and displayed a white border around the key.
 We displayed participants’ hands using the Unity MR toolkit’s default hand skin, an overlay of white and grey triangles.
-We displayed blue spheres on fingertips and grey cubes on hand joints (Figure 1 right).
-The keyboard was 44 cm 16 cm and rendered slightly transparent blue with white key labels.
+We displayed blue spheres on index fingertips and grey cubes at joints.
+The keyboard size (including four character rows plus the spacebar) was 33.0 cm 11.0 cm.
+Each letter key was 2.2 cm 2.2 cm.
+The keyboard was similar in size to a physical desktop keyboard.
+It is also similar to the virtual key sizes reported in past work, e.g.
+2.2 cm (Dudley et al., 2018), 2.5 cm (Dudley et al., 2019), and 1.9 cm (Richardson et al., 2024).
+The keyboard was rendered slightly transparent blue with white key labels.
 We used a full QWERTY keyboard including numbers, symbols, shift keys, caps lock, and backspace.
@@
 Participants used a HoloLens 2 headset while seated at a small table (Figure 1 left).
-We told participants they could type using the index fingers on one or both hands.
-During the calibration phase, participants used only one hand to avoid potential depth measurement issues caused by sensing the non-typing hand.
-During the practice and evaluation sentences, we reminded users they could type with both hands.
-We allowed participants the freedom to type with one or both hands as it better simulates real-world keyboard typing where users have the same freedom.
+The table was 91 cm wide and 51 cm deep and we placed a QR code at its top left corner.
+At the start of each condition, participants had to look at the QR code.
+This provided a spatial anchor to allow consistent placement of the virtual keyboard in each condition.
+We implemented our spatial anchoring using a MR toolkit extension (see our source code for details).
 The study lasted approximately one hour.
-Participants completed an initial questionnaire before starting the conditions.
+Participants completed an initial questionnaire with demographic questions, questions about their text entry experience, and questions about their experience with VR/AR.
 Between conditions, participants took a two-minute break.
 During the break, we instructed participants to lift the HoloLens visor to reduce eyestrain.
-They then completed a questionnaire about the previous condition.
-After all conditions, participants completed a final questionnaire.
-On the questionnaires, participants ranked various statements on a 7-point Likert scale (1=strongly disagree, 7=strongly agree).
-We also asked for open responses to various questions.
-For our exact questionnaires, see our supplementary materials.111https://osf.io/q4n72 Participants received a $15 USD gift card.
+Participants then completed a questionnaire about the previous condition.
+Participants ranked on a 7-point Likert scale (1=strongly disagree, 7=strongly agree) whether they felt the condition allowed them to enter text 1) quickly, 2) accurately, and 3) physically easily.
+We also solicited freeform feedback on what they liked and disliked about the previous condition.
+After all conditions, participants completed a final questionnaire in which they ranked conditions from best to worst and explained their ranking.
+We also asked them what they thought needed improving in AR in general.
+For our exact questionnaires, see our supplementary materials.222https://osf.io/q4n72 Participants received a $15 USD gift card.
 The order of conditions was completely counterbalanced between participants.
-Prior to each condition, participants typed a calibration sentence to ensure the keyboard’s consistent placement relative to the surface.
+Prior to each condition, participants typed a calibration sentence to set the keyboard’s height above the surface.
 If the keyboard was too far in front of the surface, the finger could actuate a key without tactile feedback.
@@
 While the calibration sentence only affected the Wall and Table conditions, we retained it in the Midair condition for consistency.
+During calibration, participants used only one hand to avoid potential depth measurement issues caused by sensing the other hand.
 Following calibration, participants typed two practice sentences and twelve evaluation sentences.
@@
 These sentences were in mixed case and included punctuation and sometimes numbers.
-Eighteen participants between the ages of 18 and 26 took part in Experiment 1.
+During the practice and evaluation sentences, we reminded users they could type with both hands.
+3.2.1. Independent variables and statistical tests.
+Experiment 1 had a single independent variable, namely the location of the keyboard (midair, table, or wall).
+We tested for statistical differences using a one-way repeated measures ANOVA.
+In cases were Mauchly’s test indicated a violation of sphericity, we report Greenhouse-Geisser corrected tests.
+For significant omnibus tests, we test for differences between conditions via Bonferroni corrected post-hoc tests.
+For Likert ratings, we used a non-parametric Friedman’s ANOVA test.
+For significant omnibus tests, we test for differences between conditions via a Bonferroni corrected non-parametric Wilcoxon signed-rank test.
+3.2.2. Participants and demographics.
+We recruited 18 participants aged 18–26.
 15 participants identified as male, 2 as female, and 1 as non-binary.
@@
 Eight participants reported augmented reality (AR) experience, including one participant who reported owning a HoloLens.
-The remaining seven participants reported limited AR experience, primarily with phone apps such as Pokémon Go.
-Throughout this paper, we test for statistical differences using repeated measures ANOVA.
-In cases were Mauchly’s test indicated a violation of sphericity, we report Greenhouse-Geisser corrected tests.
-For significant omnibus tests, we test for differences between conditions via Bonferroni corrected post-hoc tests.
-For Likert ratings, we used Friedman’s ANOVA.
-For significant omnibus tests, we test for differences between conditions via a Bonferroni corrected Wilcoxon signed-rank test.
+The remaining seven participants reported limited AR experience, primarily with phone apps (e.g.
+Pokémon Go333Pokémon Go is an AR game released in 2016 in which players locate, capture, train, and battle virtual creatures that appear in real-world locations.).
 3.3. Results
+3.3.1. Entry rate.
 Entry rate was measured in words-per-minute (WPM).
-We considered a word to be five characters (including space).
-We subtracted one from the number of characters to compensate for the unmeasured time finding and typing the first key (Wobbrock, 2007).
+As is common in text entry research, we considered a word to be five characters (including space) (Wobbrock, 2007).
+We subtracted one from the number of characters to compensate for the unmeasured time of finding and typing the first key (Wobbrock, 2007).
 We started timing from when the first key was pressed.
-We ended timing when the next button was pressed since this includes the perceptual, cognitive, and motor overheads associated with confirming the last character typed.
+We ended timing when the next button was pressed since this includes the perceptual, cognitive, and motor overheads associated with confirming the last character.
 Entry rates were 12.2, 9.1, and 7.8 WPM in the Midair, Table, and Wall conditions respectively (Figure 2 left).
@@
 Omnibus test | = , , | , , | , , |
-Pairwise | Midair Table ( ¡ ) | n/a | n/a |
-post-hoc | Midair Wall ( ¡ ) | ||
-Table Wall ( = ) |
+Pairwise | Midair Table, ¡ | n/a | n/a |
+post-hoc | Midair Wall, ¡ | ||
+tests | Table Wall, = |
+3.3.2. Error rate.
 We measured typing errors using character error rate (CER).
-CER refers to the number of character insertions, substitutions, and deletions required to change a participant’s final entered text to the reference text, relative to the characters in the reference and multiplied by 100.
+CER refers to the number of character insertions, substitutions, and deletions required to change the final entered text to the reference text, relative to the characters in the reference and multiplied by 100.
 Error rates were low in all conditions with a CER of 0.95%, 0.98%, and 0.65% in Midair, Table, and Wall respectively (Figure 2 center).
 This difference was not significant (Table 1 center).
-As a measure of how much effort participants were expending correcting errors, we calculated the number of backspaces per final output character for each phrase.
-While the final character error rates in the conditions were relatively low, we observed frequent use of backspace.
-The number of backspaces pressed per output character was 0.12 for Midair, 0.16 for Table, and 0.13 for Wall.
-While there did seem to be more backspacing in the two surface co-located keyboards (in particular for Table), this difference was not significant (Table 1 right).
-Three box plots showing the entry rate, character error rate, and backspaces per character in Experiment 1. Each plot shows the results for each of the three conditions.
-Scatter plot showing how each participant did in each condition of Experiment 1 with respect to the character error rate (x-axis) and entry rate (y-axis).
 We were curious how performance differed among participants.
@@
 passwords, account numbers, and URLs).
-We recorded the position of participants’ left and right index fingers 30 times per second.
-We also noted whenever tracking was lost for either finger.
-During the time periods participants were typing the evaluation phrases, we computed the percentage of a participant’s samples in which at least one finger was tracked.
-We found tracking success was 99.9%, 99.1%, and 98.6% in Midair, Table, and Wall respectively.
+3.3.3. Correction rate.
+As a measure of how much effort participants were expending correcting errors, we calculated the number of backspaces per final output character for each phrase.
+While the final character error rates in the conditions were relatively low, we observed frequent use of backspace.
+The number of backspaces pressed per output character was 0.12 for Midair, 0.16 for Table, and 0.13 for Wall.
+While there did seem to be more backspacing in the two surface keyboards (in particular for Table), this difference was not significant (Table 1 right).
+Three box plots showing the entry rate, character error rate, and backspaces per character in Experiment 1. Each plot shows the results for each of the three conditions.
+Scatter plot showing how each participant did in each condition of Experiment 1 with respect to the character error rate (x-axis) and entry rate (y-axis).
+3.3.4. Hand-tracking success.
+We recorded participants’ left and right index finger positions 30 times per second, including noting whether tracking was lost for either finger.
+We computed the percentage of samples where at least one finger was tracked.
+Participants’ tracking success percentages during the evaluation phrase were 99.9%, 99.1%, and 98.6% in Midair, Table, and Wall respectively.
 While this difference was significant ( = , , ), post-hoc pairwise comparisons were not significant: Midair Table ( = ), Midair Wall ( = ), and Table Wall ( = ).
-We note that just because we had a tracked location of the index fingers, this does not mean that location was necessarily accurate.
-Subjective feedback from many participants indicated hand-tracking was less accurate in Midair and Table.
+We note that just because we tracked both fingers, this does not mean the returned location was necessarily accurate.
+As we will discuss, subjective feedback from many participants indicated hand-tracking was less accurate in Midair and Table.
 We conjecture this could have resulted from a number of factors such as the proximity of the hands to the surface, occlusion of the fingers or hands between themselves, or changes due to the relative location and angle between the hands and the headset.
 Three figures of a QWERTY keyboard. Each letter is labeled with the percentage of attempts to type that letter that were successful.
-To investigate if participants’ accuracy was impacted by the spatial location of their desired key, we calculated the percentage of time participants successfully obtained the next letter in the target phrase.
+3.3.5. Accuracy over keyboard.
+To investigate if accuracy was impacted by the spatial location of a key, we calculated the percentage of time participants successfully obtained the next letter in the target phrase.
 We calculated this only on key presses in which, at the time of the key press, a participant’s current typing result was a strict prefix of the target.
-For example, if the target phrase was “Are you there?” and the participant had typed “Are you t” we would score the next key press based on a target of “h”.
-However if they had typed “Are you h” we would drop the next key press as we are unsure if they accidentally missed the “t” key or if they are writing an incorrect word like “here”.
-Filtering in this way allows us to be reasonably confident about a participant’s next intended key.
+For example, if the target phrase was “Are you there?” and the participant had typed “Are you t” we scored the next key press based on a target of “h”.
+However if they had typed “Are you h” we would drop the next key press as we are unsure if they missed the “t” key or if they are writing an incorrect word like “here”.
+This filtering allowed us to be reasonably confident about a participant’s next intended key.
 As shown in Figure 4, there did not seem to be a strong difference in accuracy based on spatial location in Midair or Table.
@@
 Image showing the Likert ratings for three statements in Experiment 1.
+3.4.1. Likert statements.
 After each condition, participants rated three statements on a 7-point Likert scale (Figure 5).
 Participants’ average ratings for “I entered text quickly” were: Table 4.0, Wall 2.9, and Midair 5.1.
-The difference between ratings was significant ().
+The difference was significant ().
 The only significant post-hoc test was that Midair was quicker than Wall ().
@@
 The only significant post-hoc test was that Midair was physically easier than Wall ().
-After each condition, we asked participants what they most liked and most disliked about the condition.
-Table 2 shows selected comments about each condition.
-Overall, a common like in Table and Wall was the tactile feedback afforded by the physical surface.
-However, in both conditions, a common dislike was problems related to the tracking of hands and failure to register key presses.
-Participants complained of physical strain in all conditions, though Table received the most positive feedback regarding ergonomics.
-For all participant comments, see our supplementary materials.222https://osf.io/q4n72
-Condition | Comment | |
-Midair | + | “Easier to lean arms against table.” |
-+ | “Not needing to touch the surface behind me helped me to lift my fingers up farther, reducing misinput.” | |
-+ | “My hands wanted to be there naturally and didn’t require as much movement between letters.” | |
-- | “Because my hands were in the air, it made physically more exhausting for longer period of time.” | |
-- | “There’s no tactile feedback.” | |
-- | “Without the wall or table it was difficult to estimate the depth of the keyboard.” | |
-Table | + | “I got to keep my arms on the table.” |
-+ | “It was a more natural typing posture.” | |
-+ | “Having the table under the keyboard allowed for a form of tactile response.” | |
-- | “It seemed I lost tracking a lot in this part; the 3D outlines of the keys felt disorienting.” | |
-- | “Accurate input required me to fully focus on my hands, causing back strain.” | |
-- | “I had to look further down than normal.” | |
-Wall | + | “I liked that the keyboard would type pretty much right as I hit the wall.” |
-+ | “Using this interface, I tended to forget that the keyboard wasn’t physically in front of me.” | |
-+ | “I didn’t have to pull away so much between letters. I didn’t really touch the table but touched the wall more.” | |
-- | “Awkward position, tiring.” | |
-- | “Hard to visualize where I needed to put my hands to type best. Typing felt very finicky.” | |
-- | “I felt like the keyboard strokes were more inaccurate than the other two tests.” |
-In a final questionnaire, participants were asked to rank the conditions.
+3.4.2. Condition ranking.
+We asked participants to rank the conditions and to explain their reasoning.
 Of the 18 participants, 13 preferred Midair, 5 preferred Table, and none preferred Wall.
@@
 Most participants ranked Wall as the worst and participants often mentioned physical strain: “vertical was the slowest and most physically taxing”, and “wall was neither comfortable nor accurate”.
+3.4.3. Likes and dislikes.
+We asked participants what they most liked and most disliked about the system after each condition.
+One author performed qualitative coding on these responses and another peer reviewed the codings and made minor improvements.
+For all participant comments, see our supplementary materials.444https://osf.io/q4n72
+Fatigue was mentioned by 8, 6, and 2 participants in the Wall, Midair, and Table conditions respectively.
+There was mixed feedback about the tactile feedback.
+8 participants said they liked the tactile feedback in the Wall condition and 4 liked it in the Table condition.
+For example, one participant liked that they could “touch the wall instead of nothing” and another liked the “physical sensation of touching the table.” None complained about the the presence of haptic feedback in those conditions.
+However, in the Midair condition, three people reported liking no tactile feedback and one disliked not having tactile feedback.
+We counted people who mentioned poor tracking in some form (unexpected key presses, failing to detect key presses, or poor tracking in general).
+9 people reported poor tracking in Wall, 5 reported it in Midair, and 13 people reported it in Table.
+Comments disliking tracking in the Table condition included “it seemed I lost tracking a lot in this part” and “sometimes I would hit one key to the left or right of the intended key.”
+3.4.4. AR improvements.
 Finally, we asked participants “what needs to be improved or changed about the AR headset’s hardware or software before you would be willing to use it in place of your mobile phone?” Requested changes were:
@@
 3.5. Discussion
-Our goal in Experiment 1 was to assess whether typing with tactile feedback outperformed or was preferred compared to typing in midair.
-We conjectured that the tactile feedback conditions, especially the Table condition which is similar to a standard physical keyboard, would perform better and be more preferred.
+Our goal in Experiment 1 was to assess whether a keyboard positioned on a surface outperformed or was preferred compared to typing in midair.
+We conjectured that the keyboards located on surfaces, especially the Table condition which is similar to a standard physical keyboard, would perform better and be more preferred.
 However, our data showed the Midair condition had a significantly higher entry rate than both Table and Wall.
 All conditions had low final error rates, though moderate use of backspace was observed.
-Furthermore, in the final questionnaire, thirteen of the eighteen participants preferred the Midair condition, with the second most preferred condition being the Table condition by the remaining five people.
-We observed throughout the study that the Table and Wall conditions reduced the reliability of key actuation and this was corroborated by comments made by participants.
+Furthermore, 13 of the 18 participants preferred the Midair condition, while the remaining 5 preferred the Table condition.
+Although Wall was ranked last, it was also the condition with the most people mentioning how they liked the tactile feedback.
+We observed throughout the study that the Table and Wall conditions reduced the reliability of key actuation and this was corroborated by participant feedback.
 This may explain the increased performance and preference for the Midair condition.
-We believe the poor performance in the Table condition may be due to our calibration process or the hand-tracking.
-During calibration, the participant had to type the word “calibrate”.
+Poor performance in the Table condition may be due to our calibration process or the hand-tracking.
+During calibration, the participant typed the word “calibrate”.
 We instructed participants to continue pushing the key until their finger touched the surface.
@@
 The latency of the hand-tracking may also have caused inaccuracies, Kim et al.
-(Kim et al., 2023) report a latency of 90 ms and tracking jitter of 1 mm using a HoloLens 2 headset.
+(2023) report a latency of 90 ms and tracking jitter of 1 mm using a HoloLens 2 headset.
 It might be possible to improve the keyboard’s placement above the surface with a modified or extended calibration procedure.
@@
 A different tracking system may have produced different results.
-However, this work provides valuable information about how well surface co-location works given the tracking of a commonly available MR headset.
-We also speculate the unreliable key triggering is related to limitations of the HoloLens’ egocentric cameras or its hand-tracking models.
-During the experiment, we saw participants try to compensate for problems with hand-tracking in a number of ways, such as moving closer to the keyboard or adjusting the viewing angle between the headset and the keyboard.
+However, this work provides valuable information about how well position a keyboard on a surface works given the tracking of a commodity MR headset.
+We also speculate the unreliable key triggering were related to limitations of the HoloLens’ egocentric cameras or its hand-tracking models.
+During the experiment, we saw participants try to compensate for hand-tracking issues in a number of ways, such as moving closer to the keyboard or adjusting the viewing angle between the headset and the keyboard.
 However, in some cases, such adjustments only made tracking less reliable.
@@
 Midair interaction can be tiring since users have to hold their hands up with little support (Boring et al., 2009; Hincapié-Ramos et al., 2014).
-Given this, we were surprised to find participants rated Midair the least tiring (Figure 5 bottom).
-As we expected, the Table was also comfortable, likely because people could rest their arms on the table similar to traditional keyboard typing.
-The Wall condition was reported to be the most physically demanding.
+While we positioned our midair keyboard such that participants could brace their forearms against the table edge, the like/dislike qualitative feedback suggested that the Midair condition was nearly as fatiguing as the Wall condition.
+However, we were surprised to find participants rated Midair the least tiring in the Likert responses (Figure 5 bottom).
+The Likert and like/dislike data both show that Table was the most comfortable, likely because people could rest their arms on the table similar to traditional keyboard typing.
+Both sets of data also show that the Wall condition was the most physically demanding.
 Although participants could rest their elbows on the table in this condition, it might have worked better for some participants than others.
-This condition is also the most different from traditional typing.
+This condition is also the most different from traditional typing and placed the keyboard the furthest away from the user.
 For consistency with the other conditions, participants stayed seated at the table while typing on the wall.
-It could be that wall typing would be more ergonomic while standing.
+Wall typing might be more ergonomic while standing.
 4. Experiment 2
@@
 The majority of participants also requested the ability to type with multiple fingers.
-Based on this, we decided to focus exclusively on a midair keyboard in Experiment 2.
+In Experiment 2, we focus exclusively on a midair typing.
 Our goals were to 1) explore if ten-finger typing is feasible on a midair keyboard, and 2) determine if eye-tracking could aid in eliminating accidental key presses.
@@
 The lack of tactile feedback also meant users often did not realize they had unintentionally pressed a key.
-Because of this, we worried that ten-finger typing, on its own, would not be competitive with index finger typing.
+Because of this, we worried that ten-finger typing, on its own, would not be competitive with index-finger typing.
 Increasingly, VR and AR headsets (including the HoloLens 2) have built-in eye-tracking.
@@
 Like Experiment 1, the keyboard was deterministic without auto-correct or word predictions.
-In the conditions using the eye-tracking filtering, we enabled all keys within a 2.6 cm radius of the user’s gaze location (roughly a one key radius).
+In the conditions using the eye-tracking filtering, we enabled all keys within a 2.6 cm radius of the user’s gaze location.
+This effectively enabled the key the user was looking at plus all adjacent keys (Figure 6).
 Keys outside this radius were disabled.
-Disabled keys were gray, while enabled keys were transparent (Figure 7).
+Disabled keys were gray, while enabled keys were transparent.
 When pressed, a key flashed a slightly transparent blue and a sound was played.
@@
 Keys outside of the radius around their gaze location were ignored.
+Image of a full QWERTY keyboard. All the keys have a grey background except for the H key and the keys adjacent to it which all have a transparent background.
 The first image shows the interface used in Experiment 2.
@@
 Each fingertip has a red sphere on it.
-Image of a full QWERTY keyboard. All the keys have a grey background except for the H key and the keys adjacent to it which all have a transparent background.
 A key was triggered when a participant’s finger pushed a key downward past a threshold of approximately one-third of a centimeter.
 In the event that multiple keys were pressed at the same time, only the furthest pressed key was registered as pressed.
-Following a key press, there would be a half-second cool down where the keyboard would not register any key presses to help prevent multiple keys from being triggered from a single tap gesture.
+Following a key press, there was a half-second cooldown where the keyboard would not register any key presses to help prevent multiple keys from being triggered from a single tap gesture.
 We used a QWERTY keyboard including numbers, symbols, shift keys, caps lock, and backspace.
-The keyboard was 35 cm 12 cm, with each letter key being 2.2 cm 2.2 cm.
-Red spherical colliders were attached to the user’s index fingers in Index and IndexEye, and all ten fingers in Ten and TenEye.
+The keyboard size (including four rows of characters plus spacebar) was 35 cm 12 cm.
+Each letter key was 2.2 cm 2.2 cm.
+Red spherical colliders were attached to a user’s index fingers in Index and IndexEye, and all ten fingers in Ten and TenEye.
 We opted for a visual design mimicking a laptop (Figure 7) since there was no need for the interface to appear in a single plane (as was the case to support the Wall condition in Experiment 1).
@@
 4.2. Procedure
-Participants used a HoloLens 2 headset while seated.
-The keyboard was placed midair in front of the participant.
-Participants were instructed to adjust the height of the chair such that they could comfortably reach the keys.
-Since the keyboard was intended to be used in midair only, no table was included in Experiment 2.
 The study took approximately one hour.
-Participants first completed an initial questionnaire and then proceeded to their first condition.
+Participants first completed an initial questionnaire.
+They then donned the HoloLens 2 headset and calibrated the eye tracker before proceeding to their first condition.
+The keyboard was placed midair in front of the participant with the bottom of the spacebar located 40 cm below eye level.
+In pilot testing, we determined this to be a comfortable seated typing location for most people.
+Participants were instructed to adjust the chair height and location to fine-tune their position with respect to the keyboard.
+Since the keyboard was intended to be used in midair only, no table was used.
+Participants were instructed to type using both hands.
 Between conditions, participants took a two-minute break and completed a questionnaire about the condition just finished.
-During this time, we instructed them to remove the headset and walk around the room to reduce eyestrain.
+During this time, we instructed participants to remove the headset and walk around the room to reduce eyestrain.
 After all conditions, participants completed a final questionnaire.
-The questionnaires were identical to Experiment 1 and are included in our supplementary materials.333https://osf.io/q4n72 Participants received a $15 USD gift card.
+The questionnaires were identical to Experiment 1 and are included in our supplementary materials.555https://osf.io/q4n72 Participants received a $15 USD gift card.
 The order of conditions was completely counterbalanced between participants.
@@
 These sentences were in mixed case and included punctuation and sometimes numbers.
-Participants were instructed to type using both hands.
-24 participants between the ages of 18 and 22 completed Experiment 2.
+4.2.1. Independent variables and statistical tests.
+We tested for statistical differences using a two-way repeated measures ANOVA.
+Our two independent variables were: 1) whether eye-tracking filtering was enabled (denoted Eye), and 2) whether participants typed with index fingers or all ten fingers (denoted Fingers).
+In the case of significant interaction effects, we tested for differences between conditions via Bonferroni corrected post-hoc tests.
+For Likert ratings, we used a non-parametric aligned rank transform (Wobbrock et al., 2011).
+4.2.2. Participants and demographics.
+We recruited 24 participants aged 18–22.
 Of these, 17 were male, 5 were female, and 2 preferred not to say.
@@
 4.3. Results
+4.3.1. Entry rate.
 We measured entry rate as in Experiment 1.
-Entry rates were 10.9 in Index, 9.4 in IndexEye, 7.1 in Ten, and 7.4 WPM in TenEye (Table 3 left and Figure 8 left).
-This difference was significant.
-All post-hoc pairwise comparisons between conditions were significant except between Ten and TenEye.
-Entry rate (WPM) | Error rate (CER %) | Backspaces per final character | |
-Index | 10.9 1.4 [8.5, 13.9] | 0.63 0.82 [0.00, 3.11] | 0.12 0.06 [0.01, 0.27] |
-IndexEye | 9.4 1.3 [6.9, 11.5] | 0.58 0.92 [0.00, 2.89] | 0.08 0.07 [0.02, 0.31] |
-Ten | 7.1 1.7 [3.1, 10.5] | 1.91 2.21 [0.00, 9.63] | 0.30 0.14 [0.07, 0.62] |
-TenEye | 7.4 1.9 [4.3, 11.0] | 1.27 1.54 [0.00, 6.52] | 0.16 0.11 [0.04, 0.39] |
-Omnibus test | , , | , , | , , |
-Pairwise | Index IndexEye, | Index IndexEye, | Index IndexEye, |
-post-hoc | Index Ten, | Index Ten, | Index Ten, |
-Index TenEye, | Index TenEye, | Index TenEye, | |
-IndexEye Ten, | IndexEye Ten, | IndexEye Ten, | |
-IndexEye TenEye, | IndexEye TenEye, | IndexEye TenEye, | |
-Ten TenEye, | Ten TenEye, | TenEye Ten, |
+Entry rates were 10.9 in Index, 9.4 in IndexEye, 7.1 in Ten, and 7.4 WPM in TenEye (Figure 8 left).
+There were significant main effects for both the use of eye-tracking and the number of fingers used (Table 2 left).
+As there was a significant interaction effect between the independent variables, we conducted post-hoc pairwise comparisons between conditions.
+All were significant except between Ten and TenEye.
 Three box plots showing the entry rate, character error rate, and backspaces per character in Experiment 2. Each plot shows the results for each of the four conditions.
 Scatter plot showing how each participant did in each condition in Experiment 2 with respect to the character error rate (x-axis) and entry rate (y-axis).
+Entry rate (WPM) | Error rate (CER %) | Backspaces per final character | ||
+Condition | Index | 10.9 1.4 [8.5, 13.9] | 0.63 0.82 [0.00, 3.11] | 0.12 0.06 [0.01, 0.27] |
+IndexEye | 9.4 1.3 [6.9, 11.5] | 0.58 0.92 [0.00, 2.89] | 0.08 0.07 [0.02, 0.31] | |
+Ten | 7.1 1.7 [3.1, 10.5] | 1.91 2.21 [0.00, 9.63] | 0.30 0.14 [0.07, 0.62] | |
+TenEye | 7.4 1.9 [4.3, 11.0] | 1.27 1.54 [0.00, 6.52] | 0.16 0.11 [0.04, 0.39] | |
+Independent | Eye | = , = , ¡ | = , = , = | = , = , ¡ |
+variable | Fingers | = , = , ¡ | = , = , ¡ | = , = , ¡ |
+Eye Fingers | = , = , ¡ | = , = , = | = , = , ¡ | |
+Pairwise | Index IndexEye, ¡ | n/a | Index IndexEye, = | |
+post-hoc | Index Ten, ¡ | Index Ten, ¡ | ||
+tests | Index TenEye, ¡ | Index TenEye, = | ||
+IndexEye Ten, ¡ | IndexEye Ten, ¡ | |||
+IndexEye TenEye, ¡ | IndexEye TenEye, ¡ | |||
+Ten TenEye, = | TenEye Ten, ¡ |
+4.3.2. Error rate.
 As in Experiment 1, we measured typing errors using character error rate (CER).
-The CER was 0.63% for Index, 0.58% for IndexEye, 1.91% for Ten, and 1.27% for TenEye (Table 3 center and Figure 8 center).
-While this difference was significant, the only significant post-hoc pairwise comparison was between IndexEye and Ten.
-To measure error correction activity, we calculated the number of backspaces per final output character for each phrase.
-The number of backspaces pressed per output character was 0.12 for Index, 0.08 for IndexEye, 0.30 for Ten, and 0.16 for TenEye (Table 3 right and Figure 8 right).
-This difference was significant.
-Post-hoc pairwise comparisons were significant with the exception of between Index and IndexEye, and between Index and TenEye.
-Notably, eye-tracking did significantly reduce the number of corrections required in ten-finger typing.
-This suggests that our filtering feature was successful at preventing some erroneous keystrokes.
+The CER was 0.63% for Index, 0.58% for IndexEye, 1.91% for Ten, and 1.27% for TenEye (Figure 8 center).
+There was a significant main effect on the number of fingers used but not for the use of eye-tracking (Table 2 center).
+Thus it seems that regardless of the use of eye-tracking filtering, errors increased for ten finger typing.
 As shown in Figure 9, numerous participants in Ten and TenEye had low entry rates of 4–8 WPM.
@@
 While entry rate was overall slow in the ten-finger conditions, the addition of eye-tracking did seem to shift performance towards a lower error rate.
+4.3.3. Correction rate.
+To measure error corrections, we calculated the number of backspaces per final output character for each phrase.
+The number of backspaces pressed per output character was 0.12 for Index, 0.08 for IndexEye, 0.30 for Ten, and 0.16 for TenEye (Figure 8 right).
+There were significant main effects for both the use of eye-tracking and the number of fingers used (Table 2 right).
+As there was a significant interaction effect, we conducted post-hoc pairwise comparisons between conditions.
+All were significant with the exception of between Index and IndexEye, and between Index and TenEye.
+Notably, eye-tracking did significantly reduce the number of corrections required in ten-finger typing.
+This suggests that our filtering feature was successful at preventing some erroneous keystrokes.
+4.3.4. Where participants looked.
 We measured the percentage of the time users were looking at a key in the Index and Ten conditions such that the key would have been enabled in IndexEye and TenEye.
@@
 A line graph showing the average eye gaze distance of participants from a pressed key over a time offset from the key press.
-Heatmap showing how often each participant used their left or right index finger.
-Heatmap showing how often each participant used each of their fingers.
 Figure 10 shows the average distance between the pressed key and eye gaze, before and after the key was pressed.
@@
 Despite users tending to look more closely at the pressed key in the IndexEye and TenEye conditions, likely due to the restraints imposed by our eye-tracking feature, all four conditions followed the same pattern of eye movement.
-This supports our theory that users will naturally tend to look where they intend to type, and indicates potential for incorporating this implicit gaze signal to improve a keyboard’s interface or algorithms (e.g.
+This supports our theory that users will naturally look where they intend to type.
+This indicates potential for incorporating this implicit gaze signal to improve a keyboard’s interface or algorithms (e.g.
 auto-correct or word predictions).
-We calculated how often participants used each of their fingers during all conditions.
-Figure 11 shows that participants in the Index and IndexEye conditions tended to share typing between their two hands.
-For example, 11 participants in the Index condition used their left index finger 40% of the time or higher.
-As shown in Figure 12, in the ten-finger conditions, in general participants used all their fingers.
-However, most participants showed a strong preference to typing with their index fingers.
-In the Ten condition, on average participants typed with their left and right index fingers 17% and 30% of the time respectively.
+4.3.5. Finger usage.
+We analyzed how often participants used each finger to type to measure how well participants followed our instructions to type using both hands.
+In the index finger conditions, participants generally used both index fingers and preferred using their right index finger.
+In the Index condition, on average, participants used their right index finger 62% of the time and ranged between 47% and 75% across participants.
+The results were similar in the IndexEye condition except that one participant used their right index finger exclusively even though we instructed all participants to try to use both hands.
+When we exclude that participant and average the others, we found that, similar to the Index condition, the right index finger was used 62% of the time and ranged from 53% to 78% across participants.
+Boxplots showing which percentage taps were performed with each finger.
+Figure 11 shows which fingers participants used in the ten-finger conditions.
+Participants used all ten fingers at least once in these conditions with only a few exceptions (four participants in Ten and one in TenEye).
+Typing with index fingers was preferred with the right index finger being used the most frequently.
+In the Ten condition, on average, participants typed with their left and right index fingers 17% and 30% of the time respectively.
 Interestingly, in the TenEye condition, this tendency was less pronounced at 12% and 21% for the left and right index fingers respectively.
@@
 4.4. Subjective feedback
+4.4.1. Likert statements.
 After each condition, participants rated three statements on a 7-point Likert scale (Figure 12).
 Participants’ average ratings for “I entered text quickly” were: Index 4.4, IndexEye 3.9, Ten 3.1, and TenEye 3.3.
-There was a significant difference between ratings (), but none of the pairwise differences were significant.
+The difference with respect to the number of fingers used was significant ( = ¡ ).
+The difference with respect to the use of eye-tracking was not significant ( = = ).
+There was no significant interaction effect ( = = ).
 Participants’ average ratings for “I entered text accurately” were: Index 5.2, IndexEye 5.8, Ten 3.5, and TenEye 4.5.
-There was a significant difference between ratings ().
-Post-hoc tests showed Index was rated more accurate than Ten (), IndexEye was more accurate than Ten (), and IndexEye was more accurate than TenEye ().
-Other pairs were non-significant.
+The differences were significant both with respect to the number of fingers () and the use of eye-tracking ().
+There was no significant interaction effect ().
 Participants’ average ratings for “Entering text was physically easy and not tiring” were: Index 4.2, IndexEye 4.4, Ten 3.7, and TenEye 4.5.
-There was not a significant difference between ratings ().
+There was no significant difference for the number of fingers (), the use of eye-tracking (), or the interaction between them ().
+4.4.2. Condition ranking.
+We asked participants to rank the conditions and to explain their reasoning.
+The Index condition was most preferred with 16 of 24 participants ranking it first.
+Some representative comments were: “I didn’t have to worry about other fingers getting in the way”, “index fingers are both faster and more accurate for this”, and “I prefer the comfort of being able to relax my other fingers”.
+The Ten condition was never ranked as the first choice and the last choice by 15 participants.
 Image showing the Likert ratings for three statements in Experiment 2.
-In a final questionnaire, participants ranked the conditions and were asked to explain their choice.
-16 of 24 participants ranked Index first.
-Of those ranking Index the best, some representative comments were: “I didn’t have to worry about other fingers getting in the way”, “index fingers are both faster and more accurate for this”, and “I prefer the comfort of being able to relax my other fingers”.
-The second most preferred condition, by seven participants, was IndexEye.
-A representative comment from a participant preferring this condition was: “index finger tracking only was the easiest, and when combined with eye-tracking, it was also the most accurate”.
-Only one participant preferred TenEye.
-Regarding the index finger conditions, participants seemed to prefer no eye-tracking, with 17 participants ranking Index above IndexEye.
-When only using index fingers, most participants seemed to think that the eye-tracking in IndexEye was more limiting than helpful.
-In the ten-finger conditions, however, participants seemed to prefer the eye-tracking feature, with 17 participants ranking TenEye above Ten.
-Participant comments specific to each condition are listed in Table 4.
-For all user feedback, see our supplementary materials.444https://osf.io/q4n72
+4.4.3. Likes and dislikes.
+We coded participants’ likes and dislikes as in Experiment 1.
+Compared to the results of the condition ranking, these responses revealed a more nuanced result.
+10 participants said they liked that Index only allowed two fingers but 7 disliked that limitation.
+After the Ten condition, 6 people said they liked being able to use all 10 fingers and 4 liked being able to use more fingers (but not all 10 fingers).
+Some representative comments were “[index fingers] were mentally easier because there was less error with less fingers able to hit random buttons” and ”the speed of typing with unrestricted 10 fingers was fun”.
+People generally disliked how hard it was to type in the Ten condition with 15 people reporting unexpected keypresses and 12 reporting poor tracking particularly for certain fingers.
+For example, one participant disliked “my fingers getting in the way while typing”.
+Eye tracking appeared to make ten finger typing less error-prone.
+After TenEye, 11 participants liked how eye tracking helped and only one disliked the unexpected keypresses in the condition.
+Despite the great reduction in unexpected keypresses in TenEye, there were two dislikes of tracking in general and nine dislikes of tracking for certain fingers.
+For example, one participant said “middle and ring fingers don’t track properly”.
+Only one participant ranked TenEye as their top ranked choice but 17 ranked it above Ten.
+IndexEye was ranked as the first choice by seven participants (the second most popular choice after Index) and was ranked above IndexEye by 17 participants.
+After the IndexEye condition, four people disliked that they couldn’t use more fingers.
+The IndexEye and TenEye conditions each had four people who disliked being required to look at the spacebar.
+This suggests a possible interface improvement in which spacebar is excluded from the eye-tracking filtering.
+4.4.4. AR improvements.
 Finally, we asked participants “what needs to be improved or changed about the AR headset’s hardware or software before you would be willing to use it in place of your mobile phone?” Requested changes were:
@@
 Reducing the bulk of the headset (1 participant).
-Condition | Comment | |
-Index | + | “It was mentally easier because there was less error with less fingers able to hit random buttons.” |
-+ | “More of a freedom to look around and plan for the next key to tap rather than be constrained by my eyes.” | |
-+ | “The other fingers couldn’t get in the way and I didn’t have to look at the space bar to press it.” | |
-- | “The press detection feels a little unreliable.” | |
-- | “I did not like the constraint of index fingers only because sometimes I would use my other fingers out of habit.” | |
-- | “I still was looking down most of it because I don’t usually type with only my index fingers.” | |
-IndexEye | + | “The eye-tracking is really neat, also definitely made me more accurate.” |
-+ | “I was able to make the fewest mistakes in this condition.” | |
-+ | “Only using index fingers made it easier to type the right keys.” | |
-- | “Needing to look at the keys I want to type slows me down a little.” | |
-- | “The lag from the eye-tracking.” | |
-- | “It took me a bit to get used to not typing with all my fingers.” | |
-Ten | + | “The speed of typing with unrestricted 10 fingers was fun.” |
-+ | “A bit less visually cluttered.” | |
-+ | “Being able to use my thumb to press space, and pinky to press shift.” | |
-- | “I got frustrated a bit so I started using fewer fingers at times because it was easier.” | |
-- | “I often found myself hitting space or another key when I didn’t want to.” | |
-- | “Ring finger tracking inaccuracies.” | |
-TenEye | + | “Definitely more accurate, was quicker typing because of that. Most false presses weren’t registered.” |
-+ | “The flexibility of using all fingers.” | |
-+ | “The eye-tracking made it harder to mis-type letters on the keyboard.” | |
-- | “The HoloLens had trouble reading the ring and pinky fingers from most natural angles.” | |
-- | “Was difficult to type quickly at first because I was getting used to the ‘look when you type’ system.” | |
-- | “Not being able to use the space bar without looking at it.” |
 4.5. Discussion
-Our goals in Experiment 2 were to 1) examine if users could type faster in midair using all ten fingers, and 2) to see if eye-tracking could be used to reduce accidental key presses.
-However, we found users had a significantly higher entry rate in the Index and IndexEye conditions than in the Ten and TenEye conditions.
-Users also expressed an overwhelming preference for the index-finger conditions.
-Users performed poorly in the Ten condition as evidenced by lower entry rates than the two index-finger conditions and more backspaces per output character then all other conditions.
-Tactile feedback or improved tracking might have improved performance.
-Participants frequently complained in open feedback about the HoloLens’ hand-tracking.
-They also reported difficulty pressing keys using their ring fingers and frustration with a lack of responsiveness from the HoloLens.
-Our hypothesis that users generally looked at the key they were attempting to press was supported by the the eye-tracking data gathered during the study.
-We examined this by noting how often people were looking near the key that they pressed in the Index and Ten conditions, which did not use eye-tracking to disable keys.
-If we had disabled keys far from their gaze point as in IndexEye and TenEye, key presses would have been excluded 32.9% and 36.0% of the time in the Index and Ten conditions respectively.
-This suggests that participants often looked where they were typing even when they were not required to do so.
+Our goals in Experiment 2 were to 1) examine if users typed faster in midair using all ten fingers, and 2) to see if eye-tracking could reduce accidental key presses.
+We found users had a significantly higher entry rate in the Index and IndexEye conditions than in the Ten and TenEye conditions.
+Users also expressed an overwhelming preference for the index finger conditions.
+Users performed poorly in the Ten condition as evidenced by lower entry rates compared to the index finger conditions and more backspaces per output character then all other conditions.
+Participants frequently complained in open feedback about the hand-tracking.
+They also reported difficulty pressing keys using their ring fingers and frustration with a lack of responsiveness of the hand-tracking.
+Our hypothesis that users looked at the key they were attempting to press was supported by our eye-tracking data.
+We examined this by noting how often people were looking near the key they pressed in the Index and Ten conditions (conditions that did not use eye-tracking to disable keys).
+If we had disabled far away keys far as in IndexEye and TenEye, key presses would have been excluded 32.9% and 36.0% of the time in the Index and Ten conditions respectively.
+This suggests participants often looked where they were typing even when they were not required to do so.
 However, the IndexEye condition only excluded 10.4% of key presses.
@@
 This result is similar to the Ten condition and may show eye-tracking successfully filtered out unintentional key presses.
-However, it is important to note we do not know how many discarded key presses were due to unintentional presses being helpfully discarded versus key presses a participant wanted to make but that were discarded because they were not looking close enough to the key.
-There is also other evidence that eye-tracking was successful.
-Firstly, responses indicated that eye-tracking generally did not increase eyestrain or physical fatigue experienced by users (Figure 13 bottom).
-Second, the addition of eye-tracking in TenEye significantly reduced the amount of backspaces per character (Figure 8 right).
-Lastly, character error rate was reduced in TenEye to levels similar to the index finger conditions.
-Overall, our results demonstrate that eye-tracking is an effective tool that should be considered when implementing ten-finger midair typing systems.
-A common complaint during the eye-tracking conditions was needing to look at the space bar to trigger it.
-Since the space bar is large and thumbs often do not move far from it when typing with ten fingers, participants felt this unnecessarily slowed typing.
-In the future, allowing space bar presses regardless of gaze position may improve results.
+However, it is important to note we do not know how many discarded key presses were due to unintentional presses being helpfully discarded versus erroneously discarded.
+A common complaint during the eye-tracking conditions was having to look at the spacebar.
+Since the spacebar was large and thumbs often do not move far from it when typing, participants felt this unnecessarily slowed typing.
+In the future, keeping the spacebar active regardless of gaze position may improve results.
 Another common complaint was that the keys were unresponsive or difficult to trigger.
-We are uncertain if this was due to hand-tracking failures or an unnecessarily deep threshold for triggering keys.
-Reducing the distance that participants needed to push keys would have made it easier for participants to trigger individual keys, potentially improving performance.
-However, a shorter distance threshold could also increase the frequency of accidental key presses.
+We are uncertain if this was due to hand-tracking issues or the particular threshold distance we used for triggering keys.
+Reducing this distance would have made it easier to trigger keys, potentially improving performance.
+However, a shorter distance could also increase accidental key presses.
 The half-second cooldown may have also limited text entry speed in some circumstance, but given all participants in our studies typed substantially slower than 24 WPM on average (i.e.
 two characters per second), we think this did not unduly slow typing.
-Additional work could investigate if a different cooldown could improve performance.
-However, similar to reducing the trigger distance, this could increase typing errors and actually slow typing.
+Additional work could investigate if different key actuation details could help improve performance.
+There is also other evidence that eye-tracking was successful.
+Firstly, the addition of eye-tracking in TenEye significantly reduced the amount of backspaces per character (Figure 8 right).
+Secondly, participants rated the stated “I entered text accurately” significantly higher in the eye-tracking conditions.
+Overall, our results demonstrate that eye-tracking is an effective tool that should be considered when implementing ten-finger midair typing systems.
+5. Limitations and Future Work
+Our entry rates in both experiments were undoubtedly limited by the deterministic nature of our keyboard.
+It is likely users could have typed substantially faster if they only had to aim approximately within a key or so of their desired letters.
+Powerful auto-correct might even be able to sift through the many false key triggers typical of ten-finger midair typing.
+However, even in the case a MR keyboard supporting fast typing via auto-correct, it would still need a fallback mode to support users entering exactly what they want.
+This can be necessary for certain types of content, or to avoid an “auto-correct trap” in which a user’s desired and correct text is repeatedly auto “corrected” into incorrect text.
+Our work can help guide how an auto-correcting MR keyboard can best support fallback to deterministic input.
+Past work on midair keyboards have commonly oriented keyboards either completely vertically (e.g.
+(Adhikary and Vertanen, 2021b; Markussen et al., 2013)) or tilted slightly away from vertical (e.g.
+(Adhikary and Vertanen, 2021a; Dudley et al., 2018)).
+However, in most past studies, participants were standing and not seated as in our study.
+We opted instead for a 15 degree tilt from horizontal for both our studies.
+This felt more comfortable in pilot testing and more closely matches users’ real-world typing experience on desktop keyboards which often have an upward tilt.
+A similar tilt of 20 degrees (Dudley, 2025) was used in another study which explored seated typing on or near a surface (Dudley et al., 2019).
+While this tilt appeared to work well for our participants, further exploration of user performance and preference with regards to keyboard tilt could be useful.
 Both experiments consisted of a single one-hour session.
@@
 Long-term use would also allow exploring whether tuning a keyboard’s models or thresholds can improve performance for a specific user.
-Our entry rates were undoubtedly limited by the deterministic nature of our keyboard.
-It is likely users could have typed substantially faster if they only had to aim approximately within a key or so of their desired letters.
-Powerful auto-correct might even be able to sift through the many false key triggers typical of ten-finger midair typing.
-However, even in the case a MR keyboard supporting fast typing via auto-correct, it would still need a fallback mode to support users entering exactly what they want.
-This can be necessary for certain types of content, or to avoid an “auto-correct trap” in which a user’s desired and correct text is repeatedly auto “corrected” into incorrect text.
-Our work here can help guide how an auto-correcting MR keyboard can best support fallback to deterministic input.
+Our participant population was young, tech savvy, and mostly male.
+While we suspect the differences we found between conditions were driven by challenges related to midair target selection and hand-tracking, we caution results could vary for other populations.
 An interesting avenue for future research would be to use auxiliary sensors (e.g.
 the AR headset’s microphone or a smartwatch) to detect taps on a physical surface instead of, or in addition to, finger collision with the keys.
-This could improve tap detection accuracy of virtual keyboards co-located with physical surfaces.
-The tactile feedback of actual contact with a surface could mitigate the problem of unintentional key presses, improving ten-finger typing.
-In our midair keyboard experiments, we used a user’s gaze location to make a hard decision that deactivated keys more than 2.6 cm away.
+This could improve tap detection accuracy of virtual targets positioned on physical surfaces.
+In particular, the tactile feedback of actual contact with a surface could mitigate the problem of unintentional key presses in ten-finger typing.
+While we a current commodity headset struggled to track all ten fingers, we found most participant complaints mostly revolved around certain fingers such as the ring and middle finger.
+An interesting future study might explore typing with a more limited set of fingers (e.g.
+the index fingers plus thumbs and pinky fingers).
+This could allow users to use some fingers for common keys (e.g.
+thumbs for spacebar and pinky finger for backspace).
+In Experiment 2, we used gaze location to make a hard decision about which keys were active.
 This threshold was determined based on pilot testing and given the eye-tracking accuracy of our headset.
@@
 It would also be interesting to investigate improving location or tap detection accuracy via a softer, probabilistic fusion of the hand and eye-tracking signals.
-We have included in our supplementary materials555https://osf.io/q4n72 detailed participant log files from both experiments.
+We have included in our supplementary materials666https://osf.io/q4n72 detailed participant log files from both experiments.
 This includes when participants triggered individual keys as well as continuously sampled observations from the hand and eye tracker for Experiment 2.
 This data could allow exploring potential gains from incorporating such changes into future midair keyboards.
-We have also included the source code for the keyboard applications we used in both experiments.
+We have also included the source code for the keyboards we used in both experiments.
 6. Conclusion
-Experiment 1 showed that a midair keyboard had significantly higher entry rates and was preferred over keyboards co-located on a surface.
+Experiment 1 showed that a midair keyboard had significantly higher entry rates and was preferred over keyboards positioned on a surface.
 Entry rates were 12.2, 9.1, and 7.8 WPM in the Midair, Table, and Wall conditions respectively.
-Questionnaire results also showed that our midair keyboard, which was arranged such that participants could rest their arms on a table, was preferred and not particularly tiring compared to the wall and table co-located keyboards.
+Questionnaire results also showed that our midair keyboard, which was arranged such that participants could rest their arms on a table, was preferred and not particularly tiring compared to the keyboards positioned on a wall or table.
 Participant feedback was generally positive about the tactile feedback and ergonomics of typing on a virtual keyboard located on a table.
-However, at present, the hand-tracking of a current commodity headset (HoloLens 2) probably need improvement to better support accurate deterministic typing near surfaces.
-In Experiment 2, we further explored midair typing by implementing typing with ten fingers and using eye-tracking to help eliminate unintentional key presses and comparing it to index-finger typing.
-While we found eye-tracking reduced the need to correct errors, ten-finger typing in midair was still slower than typing with only the index fingers.
+However, at present, the hand-tracking of a current commodity headset (HoloLens 2) needs improvement to better support accurate deterministic interactions near surfaces.
+In Experiment 2, we further explored midair typing comparing typing with ten fingers versus typing with index fingers.
+We tested using eye-tracking to help filter unintentional key presses.
+While we found eye-tracking reduced error corrections in ten-finger typing, it was still slower than index-finger typing.
 Entry rates were 10.9, 9.4, 7.1, and 7.4 WPM in the Index, IndexEye, Ten, and TenEye conditions respectively.
-In situations where midair ten-finger typing is desired, our results suggest that the addition of eye-tracking can significantly reduce backspace use without negatively impacting either entry or final error rates.
-Questionnaire results also showed no significant negative impact of the addition of eye-tracking.
+In situations where midair ten-finger typing is desired, our results suggest eye-tracking filtering can reduce backspace use without negatively impacting users’ entry or error rates.
+Subjective feedback also showed participants felt eye-tracking helped them type more accurately.
 In summary, we found that using a commodity MR headset with no auxiliary devices or sensors, novices could type at a reasonable speed of around 11–12 WPM with a error rate below 1%.
@@
 - (1)
-- Adhikary and Vertanen (2021) Jiban Adhikary and Keith Vertanen.
-2021.
+- Adhikary and Vertanen (2021a) Jiban Adhikary and Keith Vertanen.
+2021a.
 Text Entry in Virtual Environments using Speech and a Midair Keyboard.
@@
 https://doi.org/10.1109/TVCG.2021.3067776
+- Adhikary and Vertanen (2021b) Jiban Adhikary and Keith Vertanen.
+2021b.
+Typing on Midair Virtual Keyboards: Exploring Visual Designs and Interaction Styles.
+In Human-Computer Interaction – INTERACT 2021: 18th IFIP TC 13 International Conference, Bari, Italy, August 30 – September 3, 2021, Proceedings, Part IV (Bari, Italy).
+Springer-Verlag, Berlin, Heidelberg, 132–151.
+https://doi.org/10.1007/978-3-030-85610-6_9
 - Ahn et al.
@@
 IEEE, 150–159.
+- Chiossi et al.
+(2024) Francesco Chiossi, Yassmine El Khaoudi, Changkun Ou, Ludwig Sidenmark, Abdelrahman Zaky, Tiare Feuchtner, and Sven Mayer.
+2024.
+Evaluating Typing Performance in Different Mixed Reality Manifestations using Physiological Features.
+Proceedings of the ACM on Human-Computer Interaction 8, ISS, Article 542 (Oct.
+2024), 30 pages.
+https://doi.org/10.1145/3698142
 - Dudley et al.
@@
 https://doi.org/10.1109/ISMAR.2019.00027
+- Dudley (2025) John J. Dudley. 2025. Personal communication. Email correspondence.
 - Dudley et al.
@@
 - Wobbrock (2007) Jacob O Wobbrock. 2007. Measures of text entry performance. Text entry systems: Mobility, accessibility, universality (2007), 47–74.
+- Wobbrock et al.
+(2011) Jacob O Wobbrock, Leah Findlater, Darren Gergle, and James J Higgins.
+2011.
+The aligned rank transform for nonparametric factorial analyses using only anova procedures.
+In Proceedings of the SIGCHI conference on human factors in computing systems.
+143–146.
 - Xu et al.